### v2

# Solutions to multiplicative Diophantine equations in orthotopes

Abstract.
In this article, we study the number of solutions of bounded height to certain multiplicative Diophantine equations. We provide a framework on counting the number of integer solutions of such equations in an orthotope. As a consequence of our main result, we derive an effective asymptotic formula for high power moments of certain restricted divisor functions. We additionally derive an asymptotic formula for the number of singular $2\times 2$ matrices with integer entries, improving on a previous result of the first-named author.
Key words and phrases:
multiplicative energy, multiplicative Diophantine equations, determinants, integer matrices, divisor functions
2020 Mathematics Subject Classification:
11D45 (primary), 11C20 , 11D72, 15B36
1. Introduction
1.1. Set-up
Suppose we have an $m$-variable Diophantine equation
(1.1) $\phi(x_{1},\dots,x_{m})=0.$
A popular perspective on studying this equation is to consider the arithmetic statistics of the set of solutions to (1.1). A well-studied problem of this kind concerns the behaviour, for a fixed $\mathbf{b}\in{\mathbb{R}}^{m}$, of the function $\Phi(H^{\mathbf{b}})$, defined to be the number of $m$-tuples $\mathbf{x}=(x_{1},\dots,x_{m})\in{\mathbb{Z}}^{m}$, with $|x_{i}|\leq H^{b_{i}}$ for each $i=1,\dots,m$, that satisfy (1.1). In particular, taking $\mathbf{b}=\mathbf{1}_{m}$, [^1] We denote by $\mathbf{1}_{m}$ the $\bm{\mu}\in{\mathbb{N}}^{m}$ with $\mu_{j}=1$ for all $j\leqslant m$. the problem concerns the number of integer tuple solutions to (1.1) in the orthotope $[-H,H]^{m}$.
In this work, we are concerned with a particular type of multiplicative Diophantine equation that we shall describe presently. Let $k\geqslant 2$ be an integer, suppose that $\mathbf{m}=(m_{1},\dots,m_{k})\in{\mathbb{N}}^{k}$, and for each $i\leqslant k$, suppose that $\bm{\gamma}_{i}\in{\mathbb{N}}^{m_{i}}$. We consider the statistics of equations of the form
$\prod_{j_{1}=1}^{m_{1}}x_{1,j_{1}}^{\gamma_{1,j_{1}}}=\dots=\prod_{j_{k}=1}^{m_{k}}x_{k,j_{k}}^{\gamma_{k,j_{k}}}.$
For convenience, we write $\mathbf{x}=\mathbf{x}_{1}\oplus\cdots\oplus\mathbf{x}_{k}$, where $\mathbf{x}_{i}\in{\mathbb{N}}^{m_{i}}$, and denote the set of all such $\mathbf{x}$ by ${\mathbb{N}}^{\mathbf{m}}$, and say that $\mathbf{x}$ is an $\mathbf{m}$-tuple. Similarly, we write $\bm{\gamma}=\bm{\gamma}_{1}\oplus\cdots\oplus\bm{\gamma}_{k}$ and let $\gamma_{i,j}$ denote the $j^{th}$ ordinate of $\bm{\gamma}_{i}$, and for notational convenience, we let ${\mathcal{I}}^{\mathbf{m}}$ denote the set of all indices $(i,j)$ such that $1\leqslant i\leqslant k$ and $1\leqslant j\leqslant m_{i}$. With this notation, our interest therefore lies in the distribution of points $\mathbf{x}\in{\mathbb{N}}^{\mathbf{m}}$, satisfying the multiplicative Diophantine equation [^2] Here, we write $\mathbf{y}^{{\bm{\xi}}}=y_{1}^{\xi_{1}}\cdots y_{n}^{\xi_{n}}$ for any ${\bm{\xi}},\mathbf{y}\in{\mathbb{R}}^{n}_{+}$.
(1.2) $\mathbf{x}_{1}^{\bm{\gamma}_{1}}=\cdots=\mathbf{x}_{k}^{\bm{\gamma}_{k}},$
that lie inside the $\mathbf{m}$-dimensional $H^{\mathbf{b}}$-box, defined by [^3] Here, we write $z^{{\bm{\xi}}}=(z^{\xi_{1}},\ldots,z^{\xi_{n}})\in{\mathbb{R}}^{n}$ for any ${\bm{\xi}}\in{\mathbb{R}}_{+}^{n}$ and $z\in{\mathbb{R}}_{+}$.
[…]
where $b_{i,j}>0$ for all $(i,j)\in{\mathcal{I}}^{\mathbf{m}}$. Formally, we are interested in the asymptotic behaviour of the number
$\mathsf{E}(\bm{\gamma};H^{\mathbf{b}})\coloneqq\#\{\mathbf{x}\in\mathfrak{B}^{\mathbf{m}}(H^{\mathbf{b}}):\mathbf{x}_{1}^{\bm{\gamma}_{1}}=\cdots=\mathbf{x}_{k}^{\bm{\gamma}_{k}}\},$
which can be thought of as a type of weighted multiplicative energy for the set $\mathfrak{B}^{\mathbf{m}}(H^{\mathbf{b}})$. Now, the reason why we consider the solutions in the box $\mathfrak{B}^{\mathbf{m}}(H^{\mathbf{b}})$, rather than a more arbitrary bounded subset of ${\mathbb{N}}^{\mathbf{m}}$, is due to the fact that the method we use to study $\mathsf{E}(\bm{\gamma};H^{\mathbf{b}})$ produces an asymptotic formula in the most interesting, balanced case where
(1.3) $\langle\bm{\gamma}_{1},\mathbf{b}_{1}\rangle=\cdots=\langle\bm{\gamma}_{k},\mathbf{b}_{k}\rangle=N$
for some rational number $N$. Here, and in the remainder of the article, we let $\langle\cdot,\cdot\rangle:{\mathfrak{V}}\times{\mathfrak{V}}\to{\mathbb{R}}$ denote the scalar product on ${\mathfrak{V}}={\mathbb{R}}^{m}$ or ${\mathfrak{V}}={\mathbb{R}}^{\mathbf{m}}$, whichever is appropriate.
Now, the problem of estimating $\mathsf{E}(\bm{\gamma};H^{\mathbf{b}})$ for some special values of $\bm{\gamma}$ and $\mathbf{b}$ has previously appeared in the literature, e.g. [10, 36, 27, 33]. Most of these works pertain to the case where $k=2$ and $b_{i,j}=1$ for all $(i,j)\in{\mathcal{I}}^{\mathbf{m}}$, which corresponds to counting the number of positive integer solutions $x_{i,j}\leqslant H$ to the equation
$x_{1,1}^{\gamma_{1,1}}\cdots x_{1,m_{1}}^{\gamma_{1,m_{1}}}=x_{2,1}^{\gamma_{2,1}}\cdots x_{2,m_{2}}^{\gamma_{2,m_{2}}},$
[…]
It is possible, however, to consider much more general systems of equations. For example, our main result provides a method for counting the asymptotic number of integer solutions $(z_{1},z_{2},z_{3},z_{4},z_{5},z_{6})$ to
$z_{1}z_{2}^{2}z_{3}^{3}=z_{4}z_{5}^{5}=z_{6}^{6}$
with $|z_{i}|\leq H^{b_{i}}$ for $i=1,\dots,6$, as $H\to\infty$. This problem corresponds to counting $\mathsf{E}(\bm{\gamma};H^{\mathbf{b}})$ with $\bm{\gamma}=(1,2,3)\oplus(1,5)\oplus(6)$ in our notation.
Now, we note that while our study concerns the number of positive integer $\mathbf{m}$-tuple solutions to (1.2) in the box $\mathfrak{B}^{\mathbf{m}}(H^{\mathbf{b}})$, it is clear that we may also provide similar results for integer solutions by means of a simple argument.
Our paper consists of five main sections. The main results are stated in §2, with Theorem 2.1 being the most general result of this work, and Theorem 2.2 being, under some additional conditions, the strongest result. Then, Theorems 2.3 and 2.4 demonstrate how our main results can be used to derive explicit asymptotic formulæ for the power moments of restricted divisor functions.
In §3, we outline the main tools for our demonstrations, which are presented in §4. In particular, §3.1 is dedicated to a discussion of two results of de la Bretèche [8, 9], which relate the statistics of an arithmetic function $f$ to the study of the multiple Dirichlet series associated to $f$. We then, in §3.2, develop some tools for studying certain types of multiple Dirichlet series, which are essential in our application of the work of de la Bretèche.
We provide further remarks on our results and possible future directions in §5. In particular, we present a result which, under conditions that are weaker than those of Theorem 2.2, is slightly stronger than Theorem 2.1. We additionally discuss a procedure for counting solutions to (1.2) over special subsets of ${\mathbb{N}}^{\mathbf{m}}$, and suggest some problems as topics of potential future works.
1.2. Trivial bounds
We firstly discuss an elementary upper bound for $\mathsf{E}(\bm{\gamma};H^{\mathbf{b}})$, valid for any $\bm{\gamma}$ and $\mathbf{b}$. For any $\varepsilon>0$, we have
(1.4) $\mathsf{E}(\bm{\gamma};H^{\mathbf{b}})\mathop{\>\ll\>}_{\varepsilon}H^{\min(\|\mathbf{b}_{1}\|,\dots,\|\mathbf{b}_{k}\|)+\varepsilon},$
where the implied constant depends on at most $\bm{\gamma}$ and $\varepsilon$. To prove (1.4), it suffices to fix a $t\leqslant k$ and an $\mathbf{x}_{t}\in\mathfrak{B}^{m_{t}}(H^{\mathbf{b}})$, and to count the number of representations of $\mathbf{x}_{t}^{\bm{\gamma}_{t}}$ in the form $\mathbf{x}_{i}^{\bm{\gamma}_{i}}$, where $i\neq t$, and $\mathbf{x}\in{\mathbb{N}}^{m_{i}}$. Indeed, on noting that
$\prod_{\begin{subarray}{c}i=1\\ i\neq t\end{subarray}}^{k}\#\big{\{}\mathbf{x}_{i}\in{\mathbb{N}}^{m_{i}}:\mathbf{x}_{i}^{\bm{\gamma}_{i}}=\mathbf{x}_{t}^{\bm{\gamma}_{t}}\big{\}}\leqslant\#\big{\{}\mathbf{d}\in{\mathbb{N}}^{\|\bm{\gamma}_{i}\|}:d_{1}\cdots d_{\|\bm{\gamma}_{i}\|}=\mathbf{x}_{t}^{\bm{\gamma}_{t}}\big{\}}^{k-1}\mathop{\>\ll\>}_{\varepsilon}(\mathbf{x}_{t}^{\bm{\gamma}_{t}})^{(k-1)\varepsilon},$
by virtue of a standard bound for the divisor function (e.g. (1.81) in [29]), the estimate (1.4) follows on multiplying the bounds on $\mathbf{x}_{t}$ over the $m_{t}$-dimensional $H^{\mathbf{b}_{t}}$-box $\mathfrak{B}^{m_{t}}(H^{\mathbf{b}_{t}})$. Note that, in the important case where $\mathbf{b}_{i}=\mathbf{1}_{m_{i}}$, we have $\|\mathbf{b}_{i}\|=m_{i}$.
We now discuss the lower bound for the case where $\bm{\gamma}$ satisfies (1.3) and $\alpha\mathbf{b}\in{\mathbb{N}}^{\mathbf{m}}$ for some $\alpha\in{\mathbb{N}}$, that is to say that $\mathbf{b}$ has rational ordinates. These conditions naturally appear in many number theoretical applications. Trivially, we have
$\mathsf{E}(\bm{\gamma};H^{\mathbf{b}})\geqslant H^{1/\alpha}+O(1),$
which follows from the fact that $\mathbf{x}=h^{\alpha\mathbf{b}}$ is a solution to (1.2) for all $h\leqslant H^{1/\alpha}$.
[…]
(1.5) $\langle\bm{\gamma}_{1},\bm{\eta}^{u}_{1}*\mathbf{b}_{1}\rangle=\cdots=\langle\bm{\gamma}_{k},\bm{\eta}_{k}^{u}*\mathbf{b}_{k}\rangle=N_{u}\quad\text{for all}\quad u\leqslant\mu,$
and some fixed $N_{1},\ldots,N_{\mu}\in{\mathbb{Q}}$. Note that we necessarily have $N_{1}+\cdots+N_{\mu}=N$, where $N$ is as in (1.3). Moreover, in the important case where $\mathbf{b}=\mathbf{1}_{\mathbf{m}}=\mathbf{1}_{m_{1}}\oplus\cdots\oplus\mathbf{1}_{m_{k}}$, the parameter $\mu_{\bm{\gamma},\mathbf{1}_{\mathbf{m}}}$ can be seen as the maximal possible partitions of the coordinates of $\bm{\gamma}$ to collections of subsets with equal sums. For example, when
$\bm{\gamma}=(2,3,4)\oplus(3,3,3)\oplus(1,2,6),$
we may pick
$\bm{\eta}^{1}$ $=(1,0,1)\oplus(1,1,0)\oplus(0,0,1),$
$\bm{\eta}^{2}$ $=(0,1,0)\oplus(0,0,1)\oplus(1,1,0),$
which correspond to the equations
$2+0+4$ $=3+3+0=0+0+6,$
$0+3+0$ $=0+0+3=1+2+0,$
and thus shows that $\mu_{\bm{\gamma},\mathbf{1}_{(3,3,3)}}\geqslant 2$ in this case. Moreover, as $(2,3,4)\neq(3,3,3)\neq(1,2,6)$, it is clear that $\mu_{\bm{\gamma},\mathbf{1}_{(3,3,3)}}<3$, and consequently, for this choice of $\bm{\gamma}$, we have $\mu_{\bm{\gamma},\mathbf{1}_{(3,3,3)}}=2$.
Returning to general $\bm{\gamma}$ satisfying (1.5), we note that since $\alpha\mathbf{b}\in{\mathbb{N}}$, it follows that
$\mathbf{x}=d_{1}^{\bm{\eta}^{1}*\alpha\mathbf{b}}*\cdots*d_{\mu_{\bm{\gamma},\mathbf{b}}}^{\bm{\eta}^{\mu_{\bm{\gamma},\mathbf{b}}}*\alpha\mathbf{b}}\in{\mathbb{N}}^{\mathbf{m}}$
is a solution to (1.2) for any natural numbers $d_{1},\ldots,d_{\mu_{\bm{\gamma},\mathbf{b}}}\leq H^{1/\alpha}$. Therefore, we obtain the lower bound
(1.6) $\mathsf{E}(\bm{\gamma};H^{\mathbf{b}})\geqslant\#\big{\{}d_{1}^{\bm{\eta}^{1}*\alpha\mathbf{b}}*\cdots*d_{\mu_{\bm{\gamma},\mathbf{b}}}^{\bm{\eta}^{\mu_{\bm{\gamma},\mathbf{b}}}*\alpha\mathbf{b}}:d_{1},\ldots,d_{\mu_{\bm{\gamma},\mathbf{b}}}\leqslant H^{1/\alpha}\big{\}}=H^{\mu_{\bm{\gamma},\mathbf{b}}/\alpha}+O\big{(}H^{(\mu_{\bm{\gamma},\mathbf{b}}-1)/\alpha}\big{)}.$
In fact, as we will see in §5.1, the bound (1.6) is of the correct order of magnitude for certain choices of $\bm{\gamma}$ and $\mathbf{b}$. Moreover, comparing (1.4) and (1.6), it is clear that $1\leqslant\mu_{\bm{\gamma},\mathbf{b}}\leq\alpha\|\mathbf{b}_{i}\|$ for all $i\leq k$. With respect of these bounds, the main aim of this paper is to make improvements for particular cases of $\bm{\gamma}$. In some cases of particular number-theoretic interest, we derive asymptotic formulæ for $\mathsf{E}(\bm{\gamma};H^{\mathbf{b}})$, as $H\to\infty$.
1.3. Previous results
There are three main families of Diophantine equations of the form (1.2) that have previously been studied in the literature. The first family corresponds to the case where
$\bm{\gamma}=\mathbf{1}_{m}\oplus(\ell)\quad\text{and}\quad\mathbf{b}=\mathbf{1}_{m}\oplus(1)=\mathbf{1}_{(m,1)}$
for some positive integers $m$ and $\ell$. In this case, the problem is equivalent to counting the number of positive integer solutions $x_{j}\leqslant H$ to the multiplicative Diophantine equation
(1.7) $x_{1}\dots x_{m}=x_{m+1}^{\ell}.$
Before the work of de la Bretèche, Kurlberg, and Shparlinski [10], the only known results for this problem were for the smaller cases, sometimes with additional coprimality conditions. For example, for $m=2$, Tolev [46] gave the first known results on the asymptotics of $\mathsf{E}(\mathbf{1}_{2}\oplus(\ell);H^{\mathbf{1}_{(2,1)}})$. For $m=\ell=3$, de la Bretèche [7], Fouvry [20], and Heath-Brown and Moroz [28] gave related results on the number of integer solutions of this equation, with additional coprimality conditions.
For any $m,\ell\geq 2$, it was shown by de la Bretèche, Kurlberg, and Shparlinski [10] that there exists a constant $\vartheta_{m,\ell}>0$, depending at most on $m$ and $\ell$, such that
(1.8) $\mathsf{E}(\mathbf{1}_{m}\oplus(\ell);H^{\mathbf{1}_{(m,1)}})=H^{m/\ell}Q_{m,\ell}(\log H)+O(H^{m/\ell-\vartheta_{m,\ell}})$
for some real-valued polynomial $Q_{m,\ell}$ with
$\deg Q_{m,\ell}=\binom{\ell+m-1}{m-1}-m.$
They moreover demonstrated a procedure for calculating the leading coefficient of $Q_{m,\ell}$. In particular, for $m=\ell=2$, they showed the existence of an absolute constant $C$ such that
$\mathsf{E}(\mathbf{1}_{2}\oplus(2);H^{\mathbf{1}_{(2,1)}})=\dfrac{6}{\pi^{2}}H\log H+CH+O(H^{1-\vartheta_{2}}),$
which improves the error term of Tolev [46].
The second type of multiplicative equation previously studied in the literature corresponds to the case where
$\bm{\gamma}=\mathbf{1}_{m}\oplus\mathbf{1}_{m}=\mathbf{1}_{(m,m)}\quad\text{and}\quad\mathbf{b}=\mathbf{c}\oplus\mathbf{c},$
where $m$ is a natural number and $c_{j}>0$ for all $j\leqslant m$. This problem is equivalent to counting the number of positive integer solutions to
(1.9) $x_{1}\dots x_{m}=x_{m+1}\dots x_{2m},$
with $x_{j},x_{j+m}\leq H^{c_{j}}$ for all $j\leqslant m$. For general $m$, Munsch and Shparlinski [36] considered this equation to give bounds of higher moments of $\vartheta$-functions. In particular, they showed the existence of a constant $C_{m,\mathbf{c}}>0$, depending on at most $m$ and $\mathbf{c}$, such that
$\mathsf{E}(\mathbf{1}_{(m,m)};H^{\mathbf{c}\oplus\mathbf{c}})\sim C_{m,\mathbf{c}}H^{\|\mathbf{c}\|}\log^{(m-1)^{2}}H\quad\text{as}\quad H\to\infty.$
The last family comes from the work of Mastrostefano [33], who considered the arithmetic statistics of the equation
(1.10) $x_{1}x_{2}=\cdots=x_{2k-1}x_{2k}.$
This equation corresponds to (1.2) with $\bm{\gamma}=\mathbf{b}=\mathbf{1}_{2}\oplus\cdots\oplus\mathbf{1}_{2}=\mathbf{1}_{(2,\ldots,2)}$ in our notation, and is related to the $k$-th moment of the restricted divisor function $\tau(n;H)$, defined in §1.4 below. In particular, he showed the existence of a constant $D_{k}>0$ such that
$\mathsf{E}(\mathbf{1}_{(2,\ldots,2)};H^{\mathbf{1}_{(2,\ldots,2)}})\sim D_{k}H^{2}(\log H)^{2^{k}-k-1}\quad\text{as}\quad H\to\infty.$
In the case $k=2$, the constant $D_{2}$ was explicitly evaluated by the first-named author in [1], albeit following a different approach. In this article, we will later outline a procedure for determining the constant $D_{k}$ for any positive integer $k$.
1.4. Number theoretical applications
An important application of (1.9) arises in the case where $k,m=2$, which corresponds to the equation
(1.11) $x_{1}x_{2}=x_{3}x_{4}.$
As noted by Mastrostefano [33], the number of integer solutions $1\leqslant x_{1},x_{2},x_{3},x_{4}\leqslant H$ to the equation (1.11) is precisely the multiplicative energy of the first $H$ natural numbers. Moreover, this problem additionally appears in many other number theoretical problems, to which we will now mention some examples.
Firstly, the equation (1.11) can be used to study the number of $2\times 2$ integer matrices of determinant zero with entries in a box $0\leq|x_{1}|,\>|x_{2}|,\>|x_{3}|,\>|x_{4}|\leq H$. Therefore, results on the statistics of this equation complement results of Ganguly and Guria [21], Shparlinski [44], and the first-named author [1] on the number of $2\times 2$ integer matrices with bounded entries and fixed determinant. Additionally, the same equation also appears in counting the number of $2\times 2$ singular matrices whose entries are Egyptian fractions with height at most $H$. This object was studied by the first-named author along with Kuperberg, Ostafe and Shparlinski [2].
In another direction, (1.11) is related to the second moment of the restricted divisor function
[…]
Indeed, it was shown by Ayyad, Cochrane, and Zheng [3], with respect to our notation, that
(1.12) $\sum_{n}\tau(n;H)^{2}=\mathsf{E}(\mathbf{1}_{(2,2)};H^{\mathbf{1}_{(2,2)}})=\frac{2}{\zeta(2)}H^{2}\log H+AH^{2}+O\big{(}H^{19/13}(\log H)^{7/13}\big{)},$
where
$A=\frac{12\gamma-6\zeta^{\prime}(2)/\zeta(2)-3-6\zeta(2)}{3\zeta(2)}.$
A slightly weaker version of the above has recently been proven by Mastrostefano [33] and the first-named author [1] using vastly different techniques. Note that $\tau(n;H)=0$ whenever $n>H^{2}$, and thus the left-hand side of (1.12) is well-defined.
Counting instead in terms of the Euclidean norm, Louboutin and Munsch [31] showed that
#$\#\{\mathbf{x}\in\mathfrak{S}(X):x_{1}x_{2}=x_{3}x_{4}\}=\tfrac{3}{4}X^{2}\log X+O(X^{2}),$
where $\mathfrak{S}(X)=\{\mathbf{x}\in{\mathbb{N}}^{4}:\|\mathbf{x}\|_{2}\leqslant X\}$. Moreover, for a certain class of smooth weights $w:{\mathbb{Z}}^{4}\to{\mathbb{R}}_{+}$, Heath-Brown [27] showed that
$\sum_{\begin{subarray}{c}\mathbf{x}\in{\mathbb{Z}}^{4}\\ x_{1}x_{2}=x_{3}x_{4}\end{subarray}}w\big{(}X^{-1}\mathbf{x}\big{)}=C_{1}(w)X^{2}\log X+C_{2}(w)X^{2}+O_{\varepsilon}(X^{3/2+\varepsilon})$
for some constants $C_{j}(w)$, depending at most on $w$, where $C_{1}(w)>0$ can be explicitly calculated.
Now, the function $\tau(n;H)$ first appeared in Truelsen [47], under a slightly different notation, who used this function in relation to the problem of determining the irrational numbers $\alpha$ for which the pair correlation for the fractional parts of $n^{2}\alpha$ is Poissonian. Truelsen [47] and Shparlinski [43] computed the sum of this function over a short arithmetic progression. The same function also appears in Mastrostefano [32], who used it to classify maximal size product sets of random subsets of ${\mathbb{Z}}$. In this article, we discuss a generalisation of this function, the generalised restricted divisor function
[…]
where $c_{j}>0$ for all $j\leqslant m$. In particular, we will give an asymptotic for the $k$-th moment of $\tau_{m}(n;H^{\mathbf{c}})$ in §2.2. Results in this direction clearly have many potential applications in analytic number theory. Note that, we have the relation $\tau_{2}(n;H^{\mathbf{1}_{2}})=\tau(n;H)$ to the divisor function studied by Mastrosefano. We will also briefly discuss a further generalisation of $\tau_{m}(n;H^{\mathbf{c}})$ in §5.
1.5. Notations and conventions
Firstly, we note that we adopt the convention that ${\mathbb{Z}}_{+}$ and ${\mathbb{R}}_{+}$ contain 00, while ${\mathbb{N}}$ does not. The letter $p$, with or without a subscript, will always denote a prime. Similarly, any sum or product indexed by $p$ is intended to run over the primes. Any other sum or product, unless otherwise obvious, is assumed to run over the natural numbers.
We shall use $\varepsilon$ to denote an arbitrarily small positive constant, which may vary in actual value from line to line. We will additionally make use of both the usual Landau and Vinogradov notations $O,o,\sim,\ll,\gg$ (see e.g. [17, 35]), as well as the following standard conventions:
  • #$\#{\mathcal{S}}$ denotes the cardinality of a finite set ${\mathcal{S}}$.
  • $\operatorname{span}{\mathcal{S}}$ denotes the span of the set ${\mathcal{S}}$ over ${\mathbb{R}}$.
  • ${\mathrm{vol\,}}{\mathcal{T}}$ denotes the Lebesgue volume of ${\mathcal{T}}$ in an appropriate embedded subspace.
In this article, we follow the convention that the volume of a region of dimension zero, i.e. a point, is 1. Moreover, the volume of a line segment is the length of that line segment.
We also introduce the following notations, some of which are non-standard:
  • ${\mathcal{S}}^{\mathbf{m}}={\mathcal{S}}^{m_{1}}\oplus\cdots\oplus{\mathcal{S}}^{m_{k}}$, where $\mathbf{m}\in{\mathbb{N}}^{k}$, denotes the set of $\mathbf{r}=\mathbf{r}_{1}\oplus\cdots\oplus\mathbf{r}_{k}$ with $\mathbf{r}_{i}\in{\mathcal{S}}^{m_{i}}$.
  • $r_{i,j}$ denotes the $j$-th ordinate of $\mathbf{r}_{i}$.
  • $\mathbf{r}_{1}$, $\dots$, $\mathbf{r}_{k}$ are parts of $\mathbf{r}$.
  • ${\mathcal{I}}^{\mathbf{m}}$ denotes the set of integer pairs $(i,j)$ where $1\leqslant i\leqslant k$ and $1\leqslant j\leqslant m_{i}$.
  • ${\mathcal{H}}_{\mathbf{a}}^{\mathbf{m}}$ denotes the subset of $\mathbf{s}\in{\mathbb{C}}^{\mathbf{m}}$ for which ${\mathfrak{R}}s_{i,j}>a_{i,j}$ for all $i,j\in{\mathcal{I}}^{\mathbf{m}}$.
  • ${\mathfrak{R}}\mathbf{s}$, where $\mathbf{s}\in{\mathbb{C}}^{\mathbf{m}}$ is the $\bm{\sigma}\in{\mathbb{R}}^{\mathbf{m}}$ with ${\mathfrak{R}}s_{i,j}=\sigma_{i,j}$ for all $(i,j)\in{\mathcal{I}}^{\mathbf{m}}$.
  • ${\mathfrak{I}}\mathbf{s}$ is defined similarly to ${\mathfrak{R}}\mathbf{s}$.
  • $\mathbf{1}_{\mathbf{m}}=\mathbf{1}_{m_{1}}\oplus\cdots\oplus\mathbf{1}_{m_{k}}$, where $\mathbf{1}_{n}$ denotes the $\mathbf{y}\in{\mathbb{C}}^{n}$ with $y_{j}=1$ for all $j\leqslant n$.
  • $\mathbf{0}_{\mathbf{m}}=\mathbf{0}_{m_{1}}\oplus\cdots\oplus\mathbf{0}_{m_{k}}$, where $\mathbf{0}_{n}=0\cdot\mathbf{1}_{n}$.
  • $\mathbf{y}^{\mathbf{b}}=y_{1}^{b_{1}}\cdots y_{n}^{b_{n}}$ for any $\mathbf{y},\mathbf{b}\in{\mathbb{R}}^{n}$.
  • $Y^{\mathbf{b}}=(Y^{b_{1}},\ldots,Y^{b_{n}})$ for any $Y\in{\mathbb{R}}$ and $\mathbf{b}\in{\mathbb{R}}^{n}$.
  • ${\mathcal{Y}}^{{\mathcal{X}}}$ denotes the set of sequences $(y_{\mathbf{x}})_{\mathbf{x}\in{\mathcal{X}}}$ with $y_{\mathbf{x}}\in{\mathcal{Y}}$ for all $\mathbf{x}\in{\mathcal{X}}$.
  • $\langle\mathbf{a},\mathbf{b}\rangle$ denotes the standard inner product on ${\mathbb{C}}^{\mathbf{m}}$, that is, the scalar product.
  • $\|\mathbf{a}\|$ denotes the $\ell^{1}$-norm of $\mathbf{a}\in{\mathbb{R}}^{\mathbf{m}}$.
  • ${\mathcal{L}}^{\mathbf{m}}({\mathbb{C}})$ denotes the space of linear forms $\ell:{\mathbb{C}}^{\mathbf{m}}\to{\mathbb{C}}$.
  • +${\mathcal{L}}^{\mathbf{m}}_{+}({\mathbb{C}})$ denotes the subset of $\ell\in{\mathcal{L}}^{\mathbf{m}}({\mathbb{C}})$ for which the restriction of $\ell$ to ${\mathbb{R}}_{+}^{\mathbf{m}}$ maps to ${\mathbb{R}}_{+}$.
  • =${\mathcal{E}}=\{\mathbf{e}_{i,j}\}$ denotes the standard basis for ${\mathbb{C}}^{\mathbf{m}}$ over ${\mathbb{C}}$.
  • =${\mathcal{E}}^{*}=\{\mathbf{e}_{i,j}^{*}\}$ denotes the standard dual basis in ${\mathcal{L}}^{\mathbf{m}}({\mathbb{C}})$.
We will additionally write $\mathbf{a}\succ\mathbf{b}$ for some $\mathbf{a},\mathbf{b}\in{\mathbb{R}}^{\mathbf{m}}$ if $a_{i,j}>b_{i,j}$ for all $(i,j)\in{\mathcal{I}}^{\mathbf{m}}$, and we similarly define the relations $\mathbf{a}\succeq\mathbf{b}$, $\mathbf{a}\prec\mathbf{b}$, and $\mathbf{a}\preceq\mathbf{b}$.
2. Statement of new results
2.1. The main assertions
Our aim in this paper is firstly to provide a method for determining an asymptotic for the number $\mathsf{E}(\bm{\gamma};H^{\mathbf{b}})$ where $\bm{\gamma}$ is an arbitrary positive integer tuple, and $\mathbf{b}$ an arbitrary positive rational tuple. Our results, as with the results of other authors [36, 7, 9, 10, 33], relate the study of the weighted multiplicative energy (1.2) to the study of the set
[…]
of solutions to the associated additive Diophantine equation. Our first result provides a uniform treatment of all multiplicative Diophantine equations of the form (1.2), and is as follows.
Theorem 2.1.
Fix $\mathbf{m}\in{\mathbb{N}}^{k}$, and suppose that $\bm{\gamma}\in{\mathbb{N}}^{\mathbf{m}}$ and $\mathbf{b}\in{\mathbb{R}}_{>0}^{\mathbf{m}}$. With ${\mathcal{R}}_{\bm{\gamma}}$ as above, suppose that $\mathbf{a}\succ\mathbf{0}_{\mathbf{m}}$ is such that $\langle\mathbf{a},\mathbf{r}\rangle\geqslant 1$ for all non-zero $\mathbf{r}\in{\mathcal{R}}_{\bm{\gamma}}$, and write
$\kappa_{\bm{\gamma}}(\mathbf{a})=\#{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})-\operatorname{rank}{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})\quad\text{where}\quad{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})\coloneqq\{\mathbf{r}\in{\mathcal{R}}_{\bm{\gamma}}:\langle\mathbf{a},\mathbf{r}\rangle=1\}.$
Then, there exists a polynomial $Q$, of degree not exceeding $\kappa_{\bm{\gamma}}(\mathbf{a})$, such that
(2.2) $\mathsf{E}(\bm{\gamma};H^{\mathbf{b}})=H^{\langle\mathbf{a},\mathbf{b}\rangle}Q(\log H)+o(H^{\langle\mathbf{a},\mathbf{b}\rangle-\vartheta})\quad\text{as}\quad H\to\infty$
for some $\vartheta>0$, depending at most on $\mathbf{a},\mathbf{b}$, and $\bm{\gamma}$.
The above statement is derived from a result of de la Bretèche [8, 9] pertaining to sums of arithmetic functions of many variables. Note that the polynomial $Q$ is not necessarily non-zero, and thus (2.2) is not an asymptotic formula for every choice of $\mathbf{a}$, as is to be expected. Indeed, the reason why the right-hand side of (2.2) depends on $\mathbf{a}$, while the left-hand side does not, is due to the fact that $\bm{\gamma}$ explicitly determines all of the $\mathbf{a}$ for which (2.2) yields an asymptotic formula. Indeed, supposing that $\mathbf{a}$ and $\mathbf{b}$ satisfy some easily verifiable criteria, we can not only determine exactly the degree of $Q$, but also the leading coefficient. In this direction, we have the following result.
Theorem 2.2.
Assume the hypothesis of Theorem 2.1, and moreover suppose $\mathbf{a}$ is such that
(2.3) $\sum_{\mathbf{r}\in{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})}\mathbf{r}{\mathbb{Z}}_{+}={\mathcal{R}}_{\bm{\gamma}},$
and that $\mathbf{b}\in{\mathbb{Q}}_{>0}^{\mathbf{m}}$ is such that
(2.4) $\langle\bm{\gamma}_{1},\mathbf{b}_{1}\rangle=\cdots=\langle\bm{\gamma}_{k},\mathbf{b}_{k}\rangle.$
Then the polynomial $Q$ is of degree precisely $\kappa_{\bm{\gamma}}(\mathbf{a})$, and moreover satisfies
$Q(\log H)\sim(\log H)^{\kappa_{\bm{\gamma}}(\mathbf{a})}V_{\bm{\gamma}}(\mathbf{a},\mathbf{b})\prod_{p}(1-p^{-1})^{\#{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})}\sum_{n\geqslant 0}\frac{\nu_{\bm{\gamma}}(\mathbf{a};n)}{p^{n}}\quad\text{as}\quad H\to\infty,$
where $\nu_{\bm{\gamma}}(\mathbf{a};n)=\#\big{\{}\mathbf{r}\in{\mathcal{R}}_{\bm{\gamma}}:\langle\mathbf{a},\mathbf{r}\rangle=n\big{\}}$ and
[…]
Note that the condition (2.3) essentially requires that the ordinates of $\mathbf{a}$ are sufficiently small that $\|\mathbf{a}\|\leqslant\|\mathbf{a}^{\prime}\|$ for any $\mathbf{a}^{\prime}\succ\mathbf{0}_{\mathbf{m}}$ satisfying $\langle\mathbf{a}^{\prime},\mathbf{r}\rangle\geqslant 1$ for all non-zero $\mathbf{r}\in{\mathcal{R}}_{\bm{\gamma}}$. Moreover, to avoid technical complications, we consider the most-interesting, balanced case (2.4), although the method works in greater generality. We additionally remark that the region in the definition of $V_{\bm{\gamma}}(\mathbf{a},\mathbf{b})$ is of dimension $\kappa_{\bm{\gamma}}(\mathbf{a})$, and is therefore contained in an embedding of ${\mathbb{R}}_{+}^{\kappa_{\bm{\gamma}}(\mathbf{a})}$ in ${\mathbb{R}}_{+}^{{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})}$ It is the $\kappa_{\bm{\gamma}}(\mathbf{a})$-dimensional volume of the region within this embedding to which we refer in the above result.
Now, as a way of demonstrating how the above two results can be applied in practice, we will provide some applications to certain counting problems in the following sections. Later, in §5.1, we also provide a non-explicit version of Theorem 2.2, as well as some additional interesting applications.
2.2. Applications to restricted divisor functions
The condition (2.4) of Theorem 2.2 generally comes for free in most number-theoretic applications, while the condition (2.3) can be fulfilled by constructing an appropriate $\mathbf{a}$. An example that we now discuss is the problem of determining the $k$-th moment of the restricted divisor function $\tau_{m}(n;H^{\mathbf{c}})$, defined in (1.13). We firstly note that we may write
(2.5) $\sum_{n}\tau_{m}(n;H^{\mathbf{c}})^{k}=\mathsf{E}(\bm{\gamma};H^{\mathbf{b}})\quad\text{with}\quad\bm{\gamma}=\bigoplus_{i=1}^{k}\mathbf{1}_{m}\quad\text{and}\quad\mathbf{b}=\bigoplus_{i=1}^{k}\mathbf{c},$
where the sum of interest is finite for reasons similar to that in (1.13). Consequently, our problem can be restated as a generalisation of (1.9) and (1.10), where we instead wish to count the number of integer-tuple solutions $\mathbf{x}\in\mathfrak{B}^{\mathbf{m}}(H^{\mathbf{b}})$ to the multiplicative Diophantine equation
$x_{1,1}\cdots x_{1,m}=\cdots=x_{k,1}\cdots x_{k,m},$
as this is precisely the number $\mathsf{E}(\bm{\gamma};H^{\mathbf{b}})$. We can therefore derive from Theorems 2.1 and 2.2 the following result.
Theorem 2.3.
Fix integers $m,k\geqslant 1$, and suppose that $\mathbf{c}\in{\mathbb{Q}}_{>0}^{m}$. Write
$q_{m,k,\mathbf{c}}=V_{m,k,\mathbf{c}}\prod_{p}(1-p^{-1})^{m^{k}}\sum_{n\geqslant 0}\frac{1}{p^{n}}\binom{n+m-1}{m-1}^{\!\!k},$
where, with ${\mathcal{S}}_{m,k}=\{\bm{\sigma}\in\{0,1\}^{m}\oplus\cdots\oplus\{0,1\}^{m}:\|\bm{\sigma}_{i}\|=1\>\forall i\leqslant k\}$, we have written
$V_{m,k,\mathbf{c}}={\mathrm{vol\,}}\Big{\{}(v_{\bm{\sigma}})_{\bm{\sigma}\in{\mathcal{S}}_{m,k}}\in{\mathbb{R}}_{+}^{{\mathcal{S}}_{m,k}}:\sum_{\bm{\sigma}\in{\mathcal{S}}_{m,k}}v_{\bm{\sigma}}\sigma_{i,j}=c_{j}\>\forall(i,j)\in\{1,\ldots,k\}\times\{1,\ldots,m\}\Big{\}}.$
Then we have, as $H\to\infty$, the asymptotic formula
$\sum_{n}\tau_{m}(n;H^{\mathbf{c}})^{k}=q_{m,k,\mathbf{c}}H^{\|\mathbf{c}\|}(\log H)^{m^{k}-(m-1)k-1}+H^{\|\mathbf{c}\|}P_{m,k,\mathbf{c}}(\log H)+o(H^{\|\mathbf{c}\|-\vartheta_{m,k,\mathbf{c}}}),$
where $P_{m,k,\mathbf{c}}$ is a polynomial of degree not exceeding $m^{k}-(m-1)k-2$, and $\vartheta_{m,k,\mathbf{c}}>0$ is fixed, both of which depend on at most $m,k$, and $\mathbf{c}$.
The above result, in the case $m=2$ and $\mathbf{c}=\mathbf{1}_{2}$, makes explicit a result of Mastrostefano [33]. Further, taking $m=k=2$ and $\mathbf{c}=\mathbf{1}_{2}$ in the above result, we obtain a much more precise version of (1.12), and consequently improve a number of other results from the work [1] of the first-named author, which we shall discuss in §5.4. Moreover, in the case $k=2$, the above makes explicit a result of Munsch and Shparlinski [36], and in particular, if we additionally take $\mathbf{c}=\mathbf{1}_{m}$, then following an argument of Harper, Nikeghbali, and Radziwiłł [26], we see that
$V_{m,2,\mathbf{1}_{m}}=\binom{2(m-1)}{m-1}\frac{B_{m}}{m^{m-1}},$
where $B_{m}$ is the $(m-1)^{2}$-dimensional volume of the Birkhoff polytope ${\mathcal{B}}_{m}\subset{\mathbb{R}}_{+}^{m^{2}}$, for which an asymptotic formula has been established by Canfield and McKay [13]. In general, $V_{m,k,\mathbf{c}}$ represents the $m^{k}-(m-1)k-1$ dimensional volume of a polytope embedded in an $m^{k}$ dimensional space.
We now consider the generalisation of the problem (1.7) pertaining to the equation
$x_{1,1}\cdots x_{1,m}=\cdots=x_{k,1}\cdots x_{k,m}=y^{\ell}.$
Similarly to as above, this problem corresponds is to determining the average order of $\tau_{m}(n^{\ell};H^{\mathbf{c}})^{k}$, by virtue of the relation
(2.6) $\sum_{n}\tau_{m}(n^{\ell};H^{\mathbf{c}})^{k}=\mathsf{E}(\bm{\gamma};H^{\mathbf{b}})\quad\text{where}\quad\bm{\gamma}=\bigoplus_{i=1}^{k}\mathbf{1}_{m}\oplus(\ell)\quad\text{and}\quad\mathbf{b}=\bigoplus_{i=1}^{k}\mathbf{c}\oplus(\|\mathbf{c}\|/\ell).$
Hence, we have the following result as a consequence of Theorem 2.2.
Theorem 2.4.
Fix integers $m,k,\ell\geqslant 1$, and suppose that $\mathbf{c}\in{\mathbb{Q}}_{>0}^{m}$. Write
$q_{m,k,\ell,\mathbf{c}}=V_{m,k,\ell,\mathbf{c}}\prod_{p}(1-p^{-1})^{\binom{m+\ell-1}{m-1}^{k}}\sum_{n\geqslant 0}\frac{1}{p^{n}}\binom{\ell n+m-1}{m-1}^{\!\!k}$
where, with ${\mathcal{S}}_{m,k,\ell}=\{\bm{\sigma}\in\{0,1,\ldots,\ell\}^{m}\oplus\cdots\oplus\{0,1,\ldots,\ell\}^{m}:\|\bm{\sigma}_{i}\|=\ell\>\forall i\leqslant k\}$, we have written
$V_{m,k,\ell,\mathbf{c}}={\mathrm{vol\,}}\Big{\{}(v_{\bm{\sigma}})_{\bm{\sigma}\in{\mathcal{S}}_{m,k,\ell}}\in{\mathbb{R}}_{+}^{{\mathcal{S}}_{m,k,\ell}}:\sum_{\bm{\sigma}\in{\mathcal{S}}_{m,k,\ell}}v_{\bm{\sigma}}\sigma_{i,j}=c_{j}\>\forall(i,j)\in\{1,\ldots,k\}\times\{1,\ldots,m\}\Big{\}}.$
Then we have, as $H\to\infty$, the relation
$\sum_{n}\tau_{m}(n^{\ell};H^{\mathbf{c}})^{k}=q_{m,k,\ell\mathbf{c}}H^{\|\mathbf{c}\|/\ell}(\log H)^{\binom{m+\ell-1}{m-1}^{\!\!k}-(m-1)k-1}+H^{\|\mathbf{c}\|/\ell}P_{m,k,\ell,\mathbf{c}}(\log H)+o(H^{1-\vartheta_{m,k,\ell,\mathbf{c}}}),$
for some polynomial $P_{m,k,\ell,\mathbf{c}}$ of degree not exceeding $\binom{m+\ell-1}{m-1}^{\!\!k}-(m-1)k-2$, and some $\vartheta_{m,k,\ell,\mathbf{c}}>0$, both of which depend on at most $m,k,\ell$, and $\mathbf{c}$.
Taking $k=1$ and $\mathbf{c}=\mathbf{1}_{m}$ in the above result, we recover the estimate (1.8), though our expression for the area $V_{m,1,\ell,\mathbf{1}_{m}}$ is written in a slightly different manner to that given in the article [10]. Note that $V_{m,k,\ell,\mathbf{c}}$ is the $\binom{m+\ell-1}{m-1}-(m-1)k-1$ dimensional volume of a polytope embedded in a $\binom{m+\ell-1}{m-1}$ dimensional space.
3. Preliminaries
3.1. Two results of de la Bretèche
As in [10, 36, 32], our main tools will come from the theory of multiple Dirichlet series. In order to pose the problem of estimating $\mathsf{E}(\bm{\gamma};H)$ in a way that allows us to use this theory, we appeal to the work of de la Bretèche [8, 7]. Essentially, his work pertains to the asymptotic behaviour of sums of the form
$S_{f}(X^{\mathbf{b}})=\sum_{\mathbf{1}\preceq\mathbf{d}\preceq X^{\mathbf{b}}}f(\mathbf{d}),$
where $\mathbf{b}\succ\mathbf{0}_{\mathbf{m}}$ and $f:{\mathbb{N}}^{\mathbf{m}}\to{\mathbb{R}}_{+}$ are fixed. Taking $f$ to be the characteristic function of some set ${\mathcal{D}}\subset{\mathbb{N}}^{\mathbf{m}}$, we see that $S_{f}(X^{\mathbf{b}}$) simply counts the number of $\mathbf{d}\in{\mathcal{D}}$ with $d_{i,j}\leqslant X^{b_{i,j}}$ for all $(i,j)\in{\mathcal{I}}^{\mathbf{m}}$. Now, in order to properly state the aforementioned work of de la Bretèche, we must firstly develop some nomenclature. We define ${\mathcal{A}}_{f}$ to be the collection of $\mathbf{a}\in{\mathbb{R}}_{+}^{\mathbf{m}}$ such that, for any sufficiently small $\delta,\delta^{\prime}>0$, the following three properties are satisfied:
  1. (i)
    The multiple Dirichlet series
    (3.1) $D_{f}(\mathbf{s})=\sum_{\mathbf{d}\in{\mathbb{N}}^{\mathbf{m}}}f(\mathbf{d})\mathbf{d}^{-\mathbf{s}}$
    converges absolutely in the domain ${\mathcal{H}}_{\mathbf{a}}^{\mathbf{m}}$ ;
  2. (ii)
    There exists a finite family ${\mathcal{L}}_{f}(\mathbf{a})\subset{\mathcal{L}}_{+}^{\mathbf{m}}({\mathbb{C}})$ of non-zero linear forms, such that the function $G_{f}(\mathbf{a},\cdot):{\mathcal{H}}_{\mathbf{0}_{\mathbf{m}}}\to{\mathbb{C}}$, defined by the map
    $G_{f}(\mathbf{a},\cdot):\mathbf{s}\mapsto D_{f}(\mathbf{s}+\mathbf{a})\prod_{\ell\in{\mathcal{L}}_{f}(\mathbf{a})}\ell(\mathbf{s}),$
    can be analytically continued to the domain ${\mathfrak{D}}_{{\mathcal{L}}_{f}(\mathbf{a})}(0,\delta,\infty)$, where
    ${\mathfrak{D}}_{{\mathcal{L}}_{f}(\mathbf{a})}(\sigma,\delta,\Delta)=\bigcap_{\ell\in{\mathcal{L}}_{f}(\mathbf{a})}\{\mathbf{s}\in{\mathcal{H}}_{-\delta\mathbf{a}}^{\mathbf{m}}:\sigma-\delta<\langle{\mathfrak{R}}\mathbf{s},\mathbf{r}\rangle<\sigma+\Delta\}.$
  3. (iii)
    For all $\varepsilon,\sigma,\Delta>0$, there exists a constant $c_{f}(\mathbf{a};\varepsilon,\sigma,\delta,\delta^{\prime},\Delta)>0$ such that the bound
    $|G_{f}(\mathbf{a},\mathbf{s})|\leqslant c_{f}(\mathbf{a};\varepsilon,\sigma,\delta,\delta^{\prime},\Delta)(1+\|{\mathfrak{I}}\mathbf{s}\|)^{\varepsilon}\prod_{\ell\in{\mathcal{L}}_{f}(\mathbf{a})}(1+|{\mathfrak{I}}\ell(\mathbf{s})|)^{1-\delta^{\prime}\min(0,{\mathfrak{R}}\ell(\mathbf{s}))}$
    holds uniformly for $\mathbf{s}$ in the domain ${\mathfrak{D}}_{{\mathcal{L}}_{f}(\mathbf{a})}(\sigma,\delta,\Delta)$.
We note that the original result of de la Bretèche [8] is actually stronger than the following result, and is stated in a slightly different manner. Alternate versions of this result are also stated in [33, 10, 36, 9], though the form which is sufficient for our purposes is as follows.
Lemma 3.1.
Fix a $\mathbf{b}\succ\mathbf{0}_{\mathbf{m}}$ and an $f:{\mathbb{N}}^{\mathbf{m}}\to{\mathbb{R}}_{+}$ which is non-zero for infinitely many points in ${\mathbb{N}}^{\mathbf{m}}$, and let $D_{f}$, ${\mathcal{A}}_{f}$, ${\mathcal{L}}_{f}(\mathbf{a})$, and $S_{f}(X^{\mathbf{b}})$ be defined as above. Choose an $\mathbf{a}\in{\mathcal{A}}_{f}$ and write ${\mathcal{C}}_{f}(\mathbf{a})={\mathcal{L}}_{f}(\mathbf{a})\cup{\mathcal{E}}(\mathbf{a})$, where
${\mathcal{E}}(\mathbf{a})=\{\mathbf{e}_{i,j}^{*}:(i,j)\in{\mathcal{J}}(\mathbf{a})\}\quad\text{with}\quad{\mathcal{J}}(\mathbf{a})=\{(i,j)\in{\mathcal{I}}^{\mathbf{m}}:a_{i,j}=0\}.$
Then, there exists a $\vartheta>0$ and a real-valued polynomial $P$, of degree not exceeding
$\nu_{f}(\mathbf{a})=\#{\mathcal{C}}_{f}(\mathbf{a})-\operatorname{rank}{\mathcal{C}}_{f}(\mathbf{a}),$
such that
(3.2) $S_{f}(X^{\mathbf{b}})=X^{\langle\mathbf{a},\mathbf{b}\rangle}P(\log X)+o(X^{\langle\mathbf{a},\mathbf{b}\rangle-\vartheta})$
uniformly as $X\to\infty$.
We note here that the polynomial $P$ in (3.2) is not necessarily non-zero. Nonetheless, we propose in §5.1 a procedure for determining the non-triviality of $P$. Moreover, under some additional conditions, the leading coefficient of $P$ can be determined using the following result of de la Bretèche [9].
Lemma 3.2.
Assume the hypothesis of Lemma 3.1, and moreover suppose that $\ell(\mathbf{a})=1$ for all $\ell\in{\mathcal{L}}_{f}(\mathbf{a})$, and that there exist real $\beta_{\ell}>0$ such that
(3.3) $\sum_{(i,j)\in{\mathcal{I}}^{\mathbf{m}}}b_{i,j}\mathbf{e}_{i,j}^{*}=\sum_{\ell\in{\mathcal{C}}_{f}(\mathbf{a})}\beta_{\ell}\ell.$
With $G_{f}(\mathbf{a},\mathbf{s})$ as defined in condition (ii) above, suppose that there exists a function $G_{f}^{*}(\mathbf{a};\cdot):{\mathbb{C}}^{{\mathcal{C}}_{f}(\mathbf{a})}\to{\mathbb{C}}$ such that
(3.4) $G_{f}(\mathbf{a},\mathbf{s})=G_{f}^{*}(\mathbf{a};(\ell(\mathbf{s}))_{\ell\in{\mathcal{C}}_{f}(\mathbf{a})}).$
Then, we have
$P(\log X)=\frac{G_{f}(\mathbf{a},\mathbf{0}_{\mathbf{m}})\>{\mathrm{vol\,}}\Omega_{{\mathcal{L}}_{f}(\mathbf{a})}(X^{\mathbf{b}})}{X^{\langle\mathbf{a},\mathbf{b}\rangle}}+O\big{(}(\log X)^{\nu_{f}(\mathbf{a})-1}\big{)}$
where
[…]
It should be noted that the above result is a consequence of part of the statement of Théorème 2 of [8], and is incorrectly stated in the work of Mastrosefano [33] in his treatment of the system (1.10). Moreover, Munsch and Shparlinski [36] incorrectly applied Théorème 2 of [8] to their study of the system (1.9), where they should have instead utilised a result similar to Lemma 3.2 above. Nonetheless, the results in all of the aforementioned works can be recovered by appealing to Theorem 2.3 of this article.
Now, in order to be able to properly apply the above two results, we must firstly develop some knowledge on multiple Dirichlet series of the form $D_{f}(\mathbf{s})$, defined in (3.1). We shall do this in the following section.
3.2. Preliminaries on multiple Dirichlet series
Supposing that ${\mathfrak{V}}\cong{\mathbb{N}}^{n_{1}}$ and ${\mathfrak{S}}\cong{\mathbb{C}}^{n_{2}}$ for some $n_{1},n_{2}\in{\mathbb{N}}$, the theory of multiple Dirichlet series is primarily concerned with the analytic properties of series of the form
$\sum_{\mathfrak{v}\in{\mathfrak{V}}}a_{\mathfrak{v}}\exp(-\varrho_{\mathfrak{v}}(\mathbf{s})),$
where $a_{\mathfrak{v}}\in{\mathbb{C}}$ and $\varrho_{\mathfrak{v}}:{\mathfrak{S}}\to{\mathbb{C}}$ is linear for all $\mathfrak{v}\in{\mathfrak{V}}$. A complete treatment of this topic is not available in the literature, though the works [34, 40, 11, 12, 16, 19] provide a number of fundamental results pertaining to certain types of these series. In this article, we are primarily concerned with the type of multiple Dirichlet series studied by de la Bretèche [8, 7] and Onozuka [37]. In particular, we are interested in studying series of the form (3.1) in the cases where $f:{\mathbb{N}}^{\mathbf{m}}\to{\mathbb{R}}_{+}$ is polynomially bounded, that is, where there exists a ${\bm{\xi}}\in{\mathbb{N}}^{\mathbf{m}}$ and a $c>0$ for which $0\leqslant f(\mathbf{d})\leqslant c\mathbf{d}^{{\bm{\xi}}}$ for all $\mathbf{d}\in{\mathbb{N}}^{\mathbf{m}}$. Now, we are interested specifically in the case where $f$ is multiplicative, that is, in the case where, for any two distinct primes $p$ and $q$, we have [^5] Here, $\mathbf{x}*\mathbf{y}$ denotes the Hadamard product of $\mathbf{x}$ and $\mathbf{y}$.
$f(p^{\mathbf{u}}*q^{\mathbf{v}})=f(p^{\mathbf{u}})f(q^{\mathbf{v}})$
for all $\textbf{u},\textbf{v}\in{\mathbb{Z}}_{+}^{\mathbf{m}}$. Now, in the case where $f$ is multiplicative, we may express $D_{f}(\mathbf{s})$ as an Euler product of the form
(3.5) $D_{f}(\mathbf{s})=\prod_{p}D_{f}(\mathbf{s};p)\quad\text{where}\quad D_{f}(\mathbf{s};p)=\sum_{\mathbf{r}\in{\mathbb{Z}}_{+}^{\mathbf{m}}}f(p^{\mathbf{r}})p^{-\langle{\mathbf{s},\mathbf{r}\rangle}}.$
It is well-known (e.g. [4, 18, 45]) that a multiple Dirichlet series and its associated Euler product converge and diverge simultaneously. Now, it is through the Euler product representation that the absolute convergence of $D_{f}$ can be determined by studying an appropriate semi-additive function, hence the relation between the solutions of the multiplicative Diophantine equation (1.2) and the solutions to the associated additive Diophantine equation (2.1).
Now, through the Euler product representation, the study of the series $D_{f}$ can be related in some form to the study of certain $L$-functions. Indeed, the standard method of developing ratios conjectures for families of $L$-functions makes critical use of the theory of multiple Dirichlet series, e.g. in the recent work of Čech [14, 15] and Gao and Zhao [22, 23]. For applications in the analytic theory of $L$-functions, however, it is not always necessary to know precisely what the region of absolute convergence for $D_{f}$ is, only that it exists and satisfies certain properties. These properties can often be deduced from a classical result of Bochner [6, 5].
Now, of the following two results, the first describes explicitly what the region of absolute convergence of $D_{f}$ is, and the second shows that the properties (ii) and (iii) of Lemma 3.1 are satisfied for our specific choice of $f$.
Proposition 3.3.
Suppose that ${\mathcal{N}}\subseteq{\mathbb{N}}^{\mathbf{m}}$ is closed under the Hadamard product, and let ${\mathcal{R}}_{{\mathcal{N}}}$ denote the collection of $\mathbf{r}\in{\mathbb{Z}}_{+}^{\mathbf{m}}$ such that the product
$\prod_{p}\frac{p^{\sigma}-1}{p^{\sigma}-\psi_{{\mathcal{N}}}(p^{\mathbf{r}})}$
is absolutely convergent and non-zero for some $\sigma\in(0,1)$, where $\psi_{{\mathcal{N}}}$ is the characteristic function of ${\mathcal{N}}$. Suppose moreover that $\{\mathbf{0}_{\mathbf{m}}\}\subset{\mathcal{F}}\subseteq{\mathcal{R}}_{{\mathcal{N}}}$ is closed under addition, and define the multiplicative function $f:{\mathbb{N}}^{\mathbf{m}}\to\{0,1\}$ by the relation $f(p^{\mathbf{r}})=\chi_{{\mathcal{F}}}(\mathbf{r})\psi_{{\mathcal{N}}}(p^{\mathbf{r}})$ for all $\mathbf{r}\in{\mathbb{Z}}_{+}^{\mathbf{m}}$ and primes $p$, where $\chi_{{\mathcal{F}}}$ is the characteristic function of ${\mathcal{F}}$. Then, the multiple Dirichlet series $D_{f}$ is absolutely convergent precisely in the region
${\mathcal{T}}_{f}=\bigcap_{\mathbf{r}\in{\mathcal{F}}^{*}}\{\mathbf{s}\in{\mathbb{N}}^{\mathbf{m}}:\langle{\mathfrak{R}}\mathbf{s},\mathbf{r}\rangle>1\},$
where ${\mathcal{F}}^{*}$ is the set of non-zero elements of ${\mathcal{F}}$.
Proposition 3.4.
Under the hypothesis of Proposition 3.3, suppose that $\mathbf{a}\in{\mathbb{R}}_{+}^{\mathbf{m}}$ is on the boundary of ${\mathcal{T}}_{f}$, and let ${\mathcal{F}}(\mathbf{a})$ denote the set of $\mathbf{r}\in{\mathcal{F}}$ for which $\langle\mathbf{a},\mathbf{r}\rangle=1$. Then, the function $G_{f}(\mathbf{a},\cdot):{\mathcal{H}}_{\mathbf{0}_{\mathbf{m}}}^{\mathbf{m}}\to{\mathbb{C}}$, defined by the map
$G_{f}(\mathbf{a},\cdot):\mathbf{s}\mapsto D_{f}(\mathbf{s}+\mathbf{a})\prod_{\mathbf{r}\in{\mathcal{F}}(\mathbf{a})}\langle\mathbf{s},\mathbf{r}\rangle,$
can be analytically continued to a holomorphic function in the domain ${\mathcal{H}}_{-\delta\mathbf{a}}^{\mathbf{m}}$, provided that $\delta>0$ is sufficiently small.
Moreover, $G_{f}(\mathbf{a},\mathbf{0}_{\mathbf{m}})$ is non-zero, and for any $\varepsilon,\sigma,\Delta>0$, there exists a constant $c_{f}(\mathbf{a};\varepsilon,\sigma,\delta,\Delta)>0$ such that the bound
$|G_{f}(\mathbf{a},\mathbf{s})|\leqslant c_{f}(\mathbf{a};\varepsilon,\sigma,\delta,\Delta)\prod_{\mathbf{r}\in{\mathcal{F}}(\mathbf{a})}(1+|\langle{\mathfrak{I}}\mathbf{s},\mathbf{r}\rangle|)^{1-\frac{1}{3}\min(0,\langle{\mathfrak{R}}\mathbf{s},\mathbf{r}\rangle)+\varepsilon}$
holds uniformly in the domain ${\mathfrak{D}}_{{\mathcal{L}}_{f}(\mathbf{a})}(\sigma,\delta,\Delta)$, where ${\mathcal{L}}_{f}(\mathbf{a})=\{\ell:\mathbf{s}\mapsto\langle\mathbf{s},\mathbf{r}\rangle:\mathbf{r}\in{\mathcal{F}}(\mathbf{a})\}$.
It is clear that the above two results can be used, for any $f$ satisfying the hypothesis of Proposition 3.3, to find an optimal corresponding $\mathbf{a}$ that satisfies the hypotheses of Lemmata 3.1 and 3.2. Of course, in order to use Lemmata 3.1 and 3.2 to derive Theorems 2.1 and 2.2, we are only required to establish Propositions 3.3 and 3.4 in the case where ${\mathcal{N}}={\mathbb{N}}^{\mathbf{m}}$. Though, by demonstrating the above more general results, we can prove a slightly more general result, as we will discuss in §5.2.
Now, in order to establish Propositions 3.3 and 3.4, we firstly prove the following two results pertaining to the region of absolute convergence of $D_{f}$ and the positivity of $G_{f}$ on the boundary of this region.
Lemma 3.5.
Assume that the hypothesis of Proposition 3.3 holds, and suppose that $\mathbf{a}\in{\mathbb{R}}_{+}^{\mathbf{m}}$ lies in the closure of the region of absolute convergence of $D_{f}$. Write
${\mathcal{F}}(\mathbf{a})=\{\mathbf{r}\in{\mathcal{F}}:\langle\mathbf{a},\mathbf{r}\rangle=\vartheta(\mathbf{a})\}\quad\text{where}\quad\vartheta(\mathbf{a})=\inf_{\mathbf{r}\in{\mathcal{F}}^{*}}\langle\mathbf{a},\mathbf{r}\rangle.$
Then $\vartheta(\mathbf{a})\geqslant 1$, and there exists a $\delta>0$ and a complex-valued function $\phi_{f}(\mathbf{a},\mathbf{s})$, holomorphic on ${\mathcal{H}}_{(1-\delta)\mathbf{a}}^{\mathbf{m}}$, such that
$D_{f}(\mathbf{s})=\phi_{f}(\mathbf{a},\mathbf{s})\prod_{\mathbf{r}\in{\mathcal{F}}(\mathbf{a})}\zeta(\langle\mathbf{s},\mathbf{r}\rangle)\prod_{p}\frac{1-p^{-\langle\mathbf{s},\mathbf{r}\rangle}}{1-\psi_{{\mathcal{N}}}(p^{\mathbf{r}})p^{-\langle\mathbf{s},\mathbf{r}\rangle}}\quad\text{for all}\quad\mathbf{s}\in{\mathcal{H}}_{\mathbf{a}}^{\mathbf{m}}.$
Proof.
Firstly, note that there exists an $\eta(\mathbf{a})>\vartheta(\mathbf{a})$ such that $\langle{\mathfrak{R}}\mathbf{a},\mathbf{r}\rangle\geqslant\eta(\mathbf{a})$ for all $\mathbf{r}\in{\mathcal{F}}$ with $\langle{\mathfrak{R}}\mathbf{a},\mathbf{r}\rangle\neq\vartheta(\mathbf{a})$. Without loss of generality, suppose that $\langle{\mathfrak{R}}\mathbf{a},\mathbf{r}\rangle=\eta(\mathbf{a})$ for at least one $\mathbf{r}\in{\mathcal{F}}$. Thus, we see for any $\mathbf{s}\in{\mathcal{H}}_{\mathbf{a}}^{\mathbf{m}}$ that
(3.6) $D_{f}(\mathbf{s};p)=1+\sum_{\mathbf{r}\in{\mathcal{F}}(\mathbf{a})}\psi_{{\mathcal{N}}}(p^{\mathbf{r}})p^{-\langle\mathbf{s},\mathbf{r}\rangle}+O\big{(}p^{-\eta(\mathbf{a})}\big{)}.$
Now, appealing to the fact that
$\prod_{\mathbf{r}\in{\mathcal{F}}(\mathbf{a})}\big{(}1-\psi_{{\mathcal{N}}}(p^{\mathbf{r}})p^{-\langle\mathbf{s},\mathbf{r}\rangle}\big{)}=1-\sum_{\mathbf{r}\in{\mathcal{F}}(\mathbf{a})}\psi_{{\mathcal{N}}}(p^{\mathbf{r}})p^{-\langle\mathbf{s},\mathbf{r}\rangle}+O\big{(}p^{-\eta(\mathbf{a})}\big{)},$
we see from (3.6) that
(3.7) $D_{f}(\mathbf{s};p)\prod_{\mathbf{r}\in{\mathcal{F}}(\mathbf{a})}\big{(}1-\psi_{{\mathcal{N}}}(p^{\mathbf{r}})p^{-\langle\mathbf{s},\mathbf{r}\rangle}\big{)}=1+R_{f}(\mathbf{a},\mathbf{s};p),$
where $R_{f}(\mathbf{a},\mathbf{s};p)\ll p^{-\eta(\mathbf{a})}$ for any $\mathbf{s}\in{\mathcal{H}}_{\mathbf{a}}^{\mathbf{m}}$, with the implied constant depending on $f$ and $\mathbf{a}$ alone. Now, we note that the product
(3.8) $\prod_{p\leqslant X}\big{(}1+R_{f}(\mathbf{a},\mathbf{s};p)\big{)}$
is absolutely convergent to some complex number $\phi_{f}(\mathbf{a},\mathbf{s})$ as $X\to\infty$, provided that $\eta(\mathbf{a})>1$. Consequently, taking $\delta>0$ to be sufficiently small that $\vartheta(\mathbf{a})<(1-\delta)\eta(\mathbf{a})$, we see that (3.8) is absolutely convergent in the region ${\mathcal{H}}_{(1-\delta)\mathbf{a}}^{\mathbf{m}}$, and thus therein defines a holomorphic function $\phi_{f}(\mathbf{s},\mathbf{a})$. Moreover, we note that the product
$\prod_{p\leqslant X}\prod_{\mathbf{r}\in{\mathcal{F}}(\mathbf{a})}\frac{1}{1-\psi_{{\mathcal{N}}}(p^{\mathbf{r}})p^{-\langle\mathbf{s},\mathbf{r}\rangle}}$
converges absolutely to
$\prod_{\mathbf{r}\in{\mathcal{F}}(\mathbf{a})}\zeta(\langle\mathbf{s},\mathbf{r}\rangle)\prod_{p}\frac{1-p^{-\langle\mathbf{s},\mathbf{r}\rangle}}{1-\psi_{{\mathcal{N}}}(p^{\mathbf{r}})p^{-\langle\mathbf{s},\mathbf{r}\rangle}}$
[…]
Lemma 3.6.
Under the hypothesis of Lemma 3.5, $\phi_{f}(\mathbf{a},\mathbf{a})$ is non-zero.
Proof.
Firstly note that, if $\mathbf{r}$ can be expressed as a linear combination of elements of ${\mathcal{F}}(\mathbf{a})$, then it surely must be an element of ${\mathcal{F}}$. Conversely, for any $\mathbf{r}\in{\mathcal{F}}$, we trivially have
#$\#\Big{\{}(w_{\mathbf{x}})_{\mathbf{x}\in{\mathcal{F}}(\mathbf{a})}\in\{0,1\}^{{\mathcal{F}}(\mathbf{a})}:\sum_{\mathbf{x}\in{\mathcal{F}}(\mathbf{a})}w_{\mathbf{x}}\mathbf{x}=\mathbf{r}\Big{\}}\leqslant 2^{\#{\mathcal{F}}(\mathbf{a})},$
and thus, for any sequence $(p_{\mathbf{r}})_{\mathbf{r}\in{\mathcal{F}}(\mathbf{a})}\in{\mathbb{P}}^{{\mathcal{F}}(\mathbf{a})}$, we must have
#$\#\Big{\{}(q_{\mathbf{r}})_{\mathbf{r}\in{\mathcal{F}}(\mathbf{a})}\in{\mathbb{P}}^{{\mathcal{F}}(\mathbf{a})}:\prod_{\mathbf{r}\in{\mathcal{F}}(\mathbf{a})}q_{\mathbf{r}}^{\mathbf{r}}=\prod_{\mathbf{r}\in{\mathcal{F}}(\mathbf{a})}p_{\mathbf{r}}^{\mathbf{r}}\Big{\}}\leqslant 2^{\#{\mathcal{F}}(\mathbf{a})^{2}}.$
Consequently, restricting the sum over $\mathbf{d}\in{\mathbb{N}}^{\mathbf{m}}$ in the definition of $D_{f}(\mathbf{s})$ to certain prime-power tuples, we have by positivity the bound
(3.9) $D_{f}({\mathfrak{R}}\mathbf{s})$ $\geqslant\frac{1}{2^{\#{\mathcal{F}}(\mathbf{a})^{2}}}\prod_{\mathbf{r}\in{\mathcal{F}}(\mathbf{a})}\sum_{p}\psi_{{\mathcal{N}}}(p^{\mathbf{r}})p^{-\langle{\mathfrak{R}}\mathbf{s},\mathbf{r}\rangle}=\frac{1}{2^{\#{\mathcal{F}}(\mathbf{a})^{2}}}\prod_{\mathbf{r}\in{\mathcal{F}}(\mathbf{a})}\zeta(\langle{\mathfrak{R}}\mathbf{s},\mathbf{r}\rangle)\xi_{\mathbf{r}}({\mathfrak{R}}\mathbf{s})$
for some functions $\xi_{\mathbf{r}}$, holomorphic and non-zero on the closure of ${\mathcal{H}}_{\mathbf{a}}^{\mathbf{m}}$, with the property that $\xi_{\mathbf{r}}({\mathfrak{R}}\mathbf{s})>0$ for all $\mathbf{s}\in{\mathcal{H}}_{\mathbf{a}}^{\mathbf{m}}$. Thus, applying Lemma 3.5 to (3.9), we have
$\phi_{f}(\mathbf{a},{\mathfrak{R}}\mathbf{s})$ $=D_{f}({\mathfrak{R}}\mathbf{s})\prod_{\mathbf{r}\in{\mathcal{F}}(\mathbf{a})}\frac{1}{\zeta(\langle{\mathfrak{R}}\mathbf{s},\mathbf{r}\rangle)}\prod_{p}\frac{1-\psi_{{\mathcal{N}}}(p^{\mathbf{r}})p^{-\langle{\mathfrak{R}}\mathbf{s},\mathbf{r}\rangle}}{1-p^{-\langle{\mathfrak{R}}\mathbf{s},\mathbf{r}\rangle}}$
$\geqslant\frac{1}{2^{\#{\mathcal{F}}(\mathbf{a})^{2}}}\prod_{\mathbf{r}\in{\mathcal{F}}(\mathbf{a})}\xi_{\mathbf{r}}({\mathfrak{R}}\mathbf{s})\prod_{p}\frac{1-\psi_{{\mathcal{N}}}(p^{\mathbf{r}})p^{-\langle{\mathfrak{R}}\mathbf{s},\mathbf{r}\rangle}}{1-p^{-\langle{\mathfrak{R}}\mathbf{s},\mathbf{r}\rangle}}$
for any $\mathbf{s}\in{\mathcal{H}}_{\mathbf{a}}^{\mathbf{m}}$, by virtue of the fact that $\zeta(s)$ is non-zero whenever ${\mathfrak{R}}s>1$. Thus, as $\phi_{f}(\mathbf{a},\mathbf{s})$ is holomorphic at $\mathbf{s}=\mathbf{a}$, we must have
$\phi_{f}(\mathbf{a},\mathbf{a})=\lim_{\mathbf{s}\to\mathbf{a}}\phi_{f}(\mathbf{a},\mathbf{s})=\lim_{\mathbf{s}\to\mathbf{a}}\phi_{f}(\mathbf{a},{\mathfrak{R}}\mathbf{s})\geqslant\frac{1}{2^{\#{\mathcal{F}}(\mathbf{a})^{2}}}\prod_{\mathbf{r}\in{\mathcal{F}}(\mathbf{a})}\big{(}\liminf_{\mathbf{s}\to\mathbf{a}}\xi_{\mathbf{r}}({\mathfrak{R}}\mathbf{s})\big{)}\prod_{p}\frac{p-\psi_{{\mathcal{N}}}(p^{\mathbf{r}})}{p-1}>0,$
[…]
We now have the tools sufficient for the establishment of Propositions 3.3 and 3.4.
Proof of Proposition 3.3..
Firstly note that, by virtue of Lemmata 3.5 and 3.6, $\mathbf{s}\in{\mathbb{C}}^{\mathbf{m}}$ is in the closure of the region of absolute convergence of $D_{f}$ if and only if $\vartheta({\mathfrak{R}}\mathbf{s})\geqslant 1$. Consequently, the closure of the aforementioned region is precisely
$\bigcap_{\mathbf{r}\in{\mathcal{F}}^{*}}\{\mathbf{s}\in{\mathbb{C}}^{\mathbf{m}}:\langle{\mathfrak{R}}\mathbf{s},\mathbf{r}\rangle\geqslant 1\}.$
[…]
Proof of Proposition 3.4..
Appealing to Lemma 3.5 and the fact that the $\zeta$-function has a simple pole at $s=1$ and is holomorphic elsewhere, we see that
(3.10) $G_{f}(\mathbf{a},\mathbf{s}):\mathbf{s}\mapsto\phi_{f}(\mathbf{a},\mathbf{a}+\mathbf{s})\prod_{\mathbf{r}\in{\mathcal{F}}(\mathbf{a})}\zeta(1+\langle\mathbf{s},\mathbf{r}\rangle)\langle\mathbf{s},\mathbf{r}\rangle\prod_{p}\frac{1-p^{-1-\langle\mathbf{s},\mathbf{r}\rangle}}{1-\psi_{{\mathcal{N}}}(p^{\mathbf{r}})p^{-1-\langle\mathbf{s},\mathbf{r}\rangle}}$
can be analytically continued to a holomorphic function in the domain ${\mathcal{H}}_{-\delta\mathbf{a}}^{\mathbf{m}}$, where $\delta>0$ as in Lemma 3.5. This proves the first part of the assertion. The second part of the assertion, that $G_{f}(\mathbf{a},\mathbf{0}_{\mathbf{m}})$ is non-zero, follows from the fact that $G_{f}(\mathbf{a},\mathbf{s})$ is holomorphic at $\mathbf{s}=\mathbf{0}_{\mathbf{m}}$ and therefore (3.10) gives
$G_{f}(\mathbf{a},\mathbf{0}_{\mathbf{m}})$ $=\lim_{\mathbf{s}\to\mathbf{0}_{\mathbf{m}}}\phi_{f}(\mathbf{a},\mathbf{a}+\mathbf{s})\prod_{\mathbf{r}\in{\mathcal{F}}(\mathbf{a})}\zeta(1+\langle\mathbf{s},\mathbf{r}\rangle)\langle\mathbf{s},\mathbf{r}\rangle\prod_{p}\frac{1-p^{-1-\langle\mathbf{s},\mathbf{r}\rangle}}{1-\psi_{{\mathcal{N}}}(p^{\mathbf{r}})p^{-1-\langle\mathbf{s},\mathbf{r}\rangle}}$
$=\phi_{f}(\mathbf{a},\mathbf{a})\prod_{\mathbf{r}\in{\mathcal{F}}(\mathbf{a})}\prod_{p}\frac{1-p^{-1}}{1-\psi_{{\mathcal{N}}}(p^{\mathbf{r}})p^{-1}},$
which is non-zero by virtue of Lemma 3.6.
Now, assume without loss of generality that $12\delta<$ 1. Then, appealing to the Vinogradov-Korobov estimate for the $\zeta$-function (e.g. Corollary 8.28 of [29]), we see that the bound
$\zeta(1+s)s\mathop{\>\ll\>}_{\varepsilon}(1+|{\mathfrak{I}}s|)^{1-\frac{1}{3}\min(0,{\mathfrak{R}}s)+\varepsilon}$
holds uniformly in the vertical strip $\{s\in{\mathbb{C}}:\sigma-\delta<{\mathfrak{R}}s<\sigma+\Delta\}$, where the implied constant depends on at most $\varepsilon$, $\sigma$, $\delta$, and $\Delta$. Consequently, appealing to (3.10), we have
$G_{f}(\mathbf{a},\mathbf{s})\mathop{\>\ll\>}_{\varepsilon}|\phi_{f}(\mathbf{a},\mathbf{a}+\mathbf{s})|\prod_{\mathbf{r}\in{\mathcal{F}}(\mathbf{a})}(1+|\langle{\mathfrak{I}}\mathbf{s},\mathbf{r}\rangle|)^{1-\frac{1}{3}\min(0,\langle{\mathfrak{R}}\mathbf{s},\mathbf{r}\rangle)+\varepsilon}$
[…]
3.3. The volume of certain polytopes
In this final preliminary section, we formally derive a result which is trivial in concept. The following result will be essential for our calculation of the volume $V_{\bm{\gamma}}(\mathbf{a},\mathbf{b})$ in Theorem 2.2.
Lemma 3.7.
Fix integers $n$ and $L$, with $1\leqslant L\leqslant n$, and for each $d\leqslant n$ and $\ell\leqslant L$, suppose that $\theta_{\ell,d}\geqslant 0$ are real numbers with the property that $\theta_{1,d}+\cdots+\theta_{L,d}\neq 0$ for all $d\leqslant n$. For any $\bm{\sigma}\succ\mathbf{0}_{L}$, denote by $t_{\bm{\sigma}}$ the dimension of the region
$R(\bm{\sigma})=\Big{\{}\mathbf{y}\in{\mathbb{R}}_{+}^{n}:\sum_{d\leqslant n}\theta_{\ell,d}y_{d}=\sigma_{\ell}\>\>\forall\ell\leqslant L\Big{\}}.$
Moreover, let $E_{\bm{\sigma}}\cong{\mathbb{R}}_{+}^{t_{\bm{\sigma}}}$ be such that $R(\bm{\sigma})\subset E_{\bm{\sigma}}\subset{\mathbb{R}}_{+}^{n}$, and suppose that $\bm{\sigma}_{0}\succ\mathbf{0}_{L}$ is such that $E_{\bm{\sigma}}=E_{\bm{\sigma}_{0}}$ for all $\bm{\sigma}\succ\bm{\sigma}_{0}$. Then, for any fixed $\bm{\tau}\succ\bm{\sigma}_{0}$, we have
$\limsup_{\bm{\sigma}\uparrow\bm{\tau}}|{\mathrm{vol\,}}R(\bm{\tau})-{\mathrm{vol\,}}R(\bm{\sigma})|=0.$
Proof.
Suppose that $t=t_{\bm{\sigma}_{0}}>0$, for otherwise the proof is trivial. In this case, there exists a linear transform ${\mathbb{R}}_{+}^{n}\to{\mathbb{R}}_{+}^{t}$ such that, for any $\bm{\sigma}\succ\bm{\sigma}_{0}$, we can write
$R(\bm{\sigma})=\Big{\{}\mathbf{x}\in{\mathbb{R}}_{+}^{t}:\sum_{d\leqslant t}\eta_{\ell,d}x_{d}\leqslant\sigma_{\ell}\>\>\forall\ell\leqslant L\Big{\}}$
for some $\eta_{\ell,d}\geqslant 0$, independent of $\bm{\sigma}$, with the property that $\eta_{1,d}+\cdots+\eta_{L,d}\neq 0$ for all $d\leqslant t$. Now, define
$X_{1}(\bm{\sigma})=X_{1}^{*}(\bm{\sigma})=\min_{\begin{subarray}{c}\ell\leqslant L\\ \theta_{\ell,1}\neq 0\end{subarray}}\frac{\sigma_{\ell}}{\theta_{\ell,1}}\quad\text{and}\quad\Delta_{1}(\bm{\sigma})=X_{1}(\bm{\tau})-X_{1}(\bm{\bm{\sigma}})$
so that, in the case $t=1$, we have
${\mathrm{vol\,}}R(\bm{\tau})-{\mathrm{vol\,}}R(\bm{\sigma})=\Delta_{1}(\bm{\sigma})\to 0\quad\text{as}\quad\bm{\sigma}\uparrow\bm{\tau}.$
This establishes the assertion in the case where $t=1$. To complete the proof in the case $t>1$, we follow a similar argument and so, with $2\leqslant d\leqslant t$ and $x_{1},\ldots,x_{d-1}\in{\mathbb{R}}_{+}$, write
$X_{d}(x_{1},\ldots,x_{d-1};\bm{\sigma})=\min_{\begin{subarray}{c}\ell\leqslant L\\ \theta_{\ell,d}\neq 0\end{subarray}}\frac{1}{\theta_{\ell,d}}\Big{(}\sigma_{\ell}-\sum_{u\leqslant d-1}\eta_{\ell,u}x_{u}\Big{)}\quad\text{and}\quad X_{d}^{*}(\bm{\sigma})=X_{d}(0,\ldots,0;\bm{\sigma}),$
and note that, with this nomenclature, we may write
(3.11) ${\mathrm{vol\,}}R(\bm{\sigma})=\int\limits_{0}^{X_{1}(\bm{\sigma})}\int\limits_{0}^{X_{2}(x_{1};\bm{\sigma})}\cdots\int\limits_{0}^{X_{t}(x_{1},\ldots,x_{t-1};\bm{\sigma})}\mathrm{d}x_{t}\cdots\mathrm{d}x_{1}.$
For notational convenience, we shall write
$I_{d}(x_{1},\ldots,x_{d-1};\bm{\sigma})=\int\limits_{0}^{X_{d}(x_{1},\ldots,x_{d-1};\bm{\sigma})}\cdots\int\limits_{0}^{X_{t}(x_{1},\ldots,x_{t-1};\bm{\sigma})}\mathrm{d}x_{t}\cdots\mathrm{d}x_{d}$
where $2\leqslant d\leqslant t$ and $x_{1},\ldots,x_{d-1}\in{\mathbb{R}}_{+}$, and moreover define
$I_{d}^{*}(\bm{\sigma})=\sup_{x_{1}<X_{1}(\bm{\sigma})}\sup_{x_{2}<X_{2}(x_{1};\bm{\sigma})}\cdots\sup_{x_{d-1}<X_{d-1}(x_{1},\ldots,x_{d-2};\bm{\sigma})}I_{d}(x_{1},\ldots,x_{d-1};\bm{\sigma}),$
as well as
$\Delta_{d}(\bm{\sigma})=\sup_{x_{1}<X_{1}(\bm{\sigma})}\sup_{x_{2}<X_{2}(x_{1};\bm{\sigma})}\cdots\sup_{x_{d-1}<X_{d-1}(x_{1},\ldots,x_{d-2};\bm{\sigma})}X_{d}(x_{1},\ldots,x_{d-1};\bm{\tau})-X_{d}(x_{1},\ldots,x_{d-1};\bm{\sigma}).$
Thus, on noting that each $X_{d}(x_{1},\ldots,x_{d-1};\bm{\sigma})$ is increasing as $\bm{\sigma}\uparrow\bm{\tau}$, we see from (3.11) that
${\mathrm{vol\,}}R(\bm{\tau})-{\mathrm{vol\,}}R(\bm{\sigma})$ $=\int\limits_{0}^{X_{1}(\bm{\tau})}I_{2}(x_{1};\bm{\tau})\>\mathrm{d}x_{1}-\int\limits_{0}^{X_{1}(\bm{\sigma})}I_{2}(x_{1};\bm{\sigma})\>\mathrm{d}x_{1}$
$=\int\limits_{0}^{X_{1}(\bm{\sigma})}\big{(}I_{2}(x_{1};\bm{\tau})-I_{2}(x_{1};\bm{\sigma})\big{)}\>\mathrm{d}x_{1}+\int\limits_{X_{1}(\bm{\sigma})}^{X_{1}(\bm{\tau})}I_{2}(x_{1};\bm{\tau})\>\mathrm{d}x_{1}$
(3.12) $=\int\limits_{0}^{X_{1}(\bm{\sigma})}\big{(}I_{2}(x_{1};\bm{\tau})-I_{2}(x_{1};\bm{\sigma})\big{)}\>\mathrm{d}x_{1}+O\big{(}\Delta_{1}(\bm{\sigma})I_{2}^{*}(\bm{\tau})\big{)}.$
Now, continuing this procedure, we see that we may rewrite the integral on the right-hand side of (3.3) as
$\int\limits_{0}^{X_{1}(\bm{\sigma})}\int\limits_{0}^{X_{2}(x_{1};\bm{\sigma})}\big{(}I_{3}(x_{1},x_{2};\bm{\tau})-I_{3}(x_{1},x_{2};\bm{\sigma})\big{)}\>\mathrm{d}x_{2}\>\mathrm{d}x_{1}+O\big{(}X_{1}(\bm{\sigma})\Delta_{2}(\bm{\sigma})I_{3}^{*}(\bm{\tau})\big{)}$
$\hskip 42.67912pt=\int\limits_{0}^{X_{1}(\bm{\sigma})}\int\limits_{0}^{X_{2}(x_{1};\bm{\sigma})}\cdots\int\limits_{0}^{X_{t-1}(x_{1},\ldots,x_{t-2};\bm{\sigma})}\big{(}I_{t}(x_{1},\ldots,x_{t-1};\bm{\tau})-I_{t}(x_{1},\ldots,x_{t-1};\bm{\sigma})\big{)}\>\mathrm{d}x_{t-1}\cdots\mathrm{d}x_{1}$
$\hskip 256.0748pt+O\Big{(}\sum_{3\leqslant d\leqslant t}X_{1}^{*}(\bm{\sigma})\cdots X_{d-2}^{*}(\bm{\sigma})\Delta_{d-1}(\bm{\sigma})I_{d}^{*}(\bm{\tau})\Big{)}.$
Therefore, by virtue of (3.3) and the fact that $|I_{t}(x_{1},\ldots,x_{t-1};\bm{\tau})-I_{t}(x_{1},\ldots,x_{t-1};\bm{\sigma})|\leqslant\Delta_{t}(\bm{\sigma})$, we must have
${\mathrm{vol\,}}R(\bm{\tau})-{\mathrm{vol\,}}R(\bm{\sigma})$ $\ll\Delta_{1}(\bm{\sigma})I_{2}^{*}(\bm{\tau})+X_{1}^{*}(\bm{\sigma})\cdots X_{t-1}^{*}(\bm{\sigma})\Delta_{t}(\bm{\sigma})$
$\hskip 170.71652pt+\sum_{3\leqslant d\leqslant t}X_{1}^{*}(\bm{\sigma})\cdots X_{d-2}^{*}(\bm{\sigma})\Delta_{d-1}(\bm{\sigma})I_{d}^{*}(\bm{\tau})$
$\ll\Delta_{1}(\bm{\sigma})I_{2}^{*}(\bm{\tau})+X_{1}^{*}(\bm{\tau})\cdots X_{t-1}^{*}(\bm{\tau})\Delta_{t}(\bm{\sigma})$
$\hskip 170.71652pt+\sum_{3\leqslant d\leqslant t}X_{1}^{*}(\bm{\tau})\cdots X_{d-2}^{*}(\bm{\tau})I_{d}^{*}(\bm{\tau})\Delta_{d-1}(\bm{\sigma}),$
where the implied constants are absolute. Consequently, we see that there exists a constant $C_{\bm{\tau}}>0$, depending at most on $\bm{\tau}$, $t$, and the $\eta_{\ell,d}$, such that
$|{\mathrm{vol\,}}R(\bm{\tau})-{\mathrm{vol\,}}R(\bm{\sigma})|\leqslant C_{\bm{\tau}}\big{(}\Delta_{1}(\bm{\sigma})+\cdots+\Delta_{t}(\bm{\sigma})\big{)}\to 0\quad\text{as}\quad\bm{\sigma}\uparrow\bm{\tau}.$
[…]
We now move on to the demonstrations of the main results.
4. Demonstrations
4.1. Proof of Theorem 2.1
Using Lemma 3.1 with Propositions 3.3 and 3.4, we can derive an estimation for $\mathsf{E}(\bm{\gamma};H^{\mathbf{b}})$ pertaining to arbitrary $\bm{\gamma}\in{\mathbb{N}}^{\mathbf{m}}$ and $\mathbf{b}\succ\mathbf{0}_{\mathbf{m}}$. In fact, we can more generally establish an estimate for the number
$\mathsf{E}(\bm{\gamma},{\mathcal{N}};H^{\mathbf{b}})=\#\big{\{}\mathbf{x}\in\mathfrak{B}^{\mathbf{m}}(H^{\mathbf{b}})\cap{\mathcal{N}}:\mathbf{x}_{1}^{\bm{\gamma}_{1}}=\cdots=\mathbf{x}_{k}^{\bm{\gamma}_{k}}\big{\}},$
where ${\mathcal{N}}\subseteq{\mathbb{N}}^{\mathbf{m}}$ is closed under the Hadmard product, and not too sparse. We will explain in §5.2 the advantages of considering this more general expression. In this section, we will prove a generalisation of Theorem 2.1, pertaining to the number $\mathsf{E}(\bm{\gamma},{\mathcal{N}};H^{\mathbf{b}})$, as our method easily allows for it. Indeed, we have the following proposition.
Proposition 4.1.
Suppose that $\bm{\gamma}$, $\mathbf{b}$, $\mathbf{a}$, and $\kappa_{\bm{\gamma}}(\mathbf{a})$ are as in Theorem 2.1, and that ${\mathcal{N}}\subseteq{\mathbb{N}}^{\mathbf{m}}$ is as in Proposition 3.3. Then, there exists a polynomial $P$, of degree not exceeding $\kappa_{\bm{\gamma}}(\mathbf{a})$, such that
$\mathsf{E}\big{(}\bm{\gamma},{\mathcal{N}};H^{\mathbf{b}}\big{)}=H^{\langle\mathbf{a},\mathbf{b}\rangle}P(\log H)+o(H^{\langle\mathbf{a},\mathbf{b}\rangle-\vartheta})\quad\text{as}\quad H\to\infty$
for some $\vartheta>0$, depending only on $\bm{\gamma}$ and $\mathbf{a}$ and $\mathbf{b}$.
Proof.
Let $\chi_{\bm{\gamma}}$ denote the characteristic function of the set ${\mathcal{D}}_{\bm{\gamma}}=\{\mathbf{d}\in{\mathbb{N}}^{\mathbf{m}}:\mathbf{d}_{1}^{\bm{\gamma}_{1}}=\cdots=\mathbf{d}_{k}^{\bm{\gamma}_{k}}\}$, and let $\psi_{\mathcal{N}}$ denote the characteristic function of the set ${\mathcal{N}}$. Then, we very clearly have
(4.1) $\mathsf{E}\big{(}\bm{\gamma},{\mathcal{N}};H^{\mathbf{b}}\big{)}=\#\big{(}{\mathcal{D}}_{\bm{\gamma}}\cap{\mathcal{N}}\cap\mathfrak{B}^{\mathbf{m}}(H^{\mathbf{b}})\big{)}=S_{\psi_{{\mathcal{N}}}\chi_{\bm{\gamma}}}(H^{\mathbf{b}}).$
Now, by virtue of Propositions 3.3 and 3.4, the hypothesis of Lemma 3.1 is satisfied for our choice of $f=\psi_{{\mathcal{N}}}\chi_{\bm{\gamma}}$ and ${\mathcal{L}}_{\psi_{{\mathcal{N}}}\chi_{\bm{\gamma}}}(\mathbf{a})={\mathcal{L}}_{\chi_{\bm{\gamma}}}(\mathbf{a})$, where
${\mathcal{L}}_{\chi_{\bm{\gamma}}}(\mathbf{a})=\{\ell:\mathbf{s}\mapsto\langle\mathbf{s},\mathbf{r}\rangle:\mathbf{r}\in{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})\}.$
[…]
On taking ${\mathcal{N}}={\mathbb{N}}^{\mathbf{m}}$ in the above result, we obtain Theorem 2.1. We will discuss the corresponding results for other choices of ${\mathcal{N}}$ in §5.2.
4.2. Proof of Theorem 2.2
In this section, we will prove, in two steps, an asymptotic formula for the number $\mathsf{E}(\bm{\gamma},{\mathcal{N}};H^{\mathbf{b}})$, under the hypothesis of Theorem 2.2. Firstly, we use Lemma 3.2 and Proposition 3.4 to prove the following result on the leading term of the polynomial.
Proposition 4.2.
Assume the hypothesis of Proposition 4.1, and moreover that $\mathbf{a}$ and $\mathbf{b}$ are as in Theorem 2.2. For any natural number $n$ and prime $p$, define the number
$\nu_{\bm{\gamma},{\mathcal{N}}}(\mathbf{a};n,p)=\#\{\mathbf{r}\in{\mathcal{R}}_{\bm{\gamma},{\mathcal{N}}}(p):\langle\mathbf{a},\mathbf{r}\rangle=n\}\quad\text{where}\quad{\mathcal{R}}_{\bm{\gamma},{\mathcal{N}}}(p)=\big{\{}\mathbf{r}\in{\mathcal{R}}_{\bm{\gamma}}:p^{\mathbf{r}}\in{\mathcal{N}}\>\forall(i,j)\in{\mathcal{I}}^{\mathbf{m}}\big{\}}.$
Then, the polynomial $P$ in Proposition 4.1 satisfies
$P(\log H)=\frac{{\mathrm{vol\,}}\Omega_{{\mathcal{L}}_{\chi_{\bm{\gamma}}}(\mathbf{a})}(H^{\mathbf{b}})}{H^{\langle\mathbf{a},\mathbf{b}\rangle}}\prod_{p}(1-p^{-1})^{\#{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})}\sum_{n\geqslant 0}\frac{\nu_{\bm{\gamma},{\mathcal{N}}}(\mathbf{a};n,p)}{p^{n}}+O\big{(}(\log H)^{\kappa_{\bm{\gamma}}(\mathbf{a})-1}\big{)}$
as $H\to\infty$, where $\Omega_{{\mathcal{L}}_{\chi_{\bm{\gamma}}}(\mathbf{a})}(H^{\mathbf{b}})$ is as in Lemma 3.2, with ${\mathcal{L}}_{\chi_{\bm{\gamma}}}(\mathbf{a})=\{\ell:\mathbf{s}\to\langle\mathbf{s},\mathbf{r}\rangle:\mathbf{r}\in{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})\}$.
Proof.
We firstly note that the additional conditions imply that any $\mathbf{r}\in{\mathcal{R}}_{\bm{\gamma}}$ can be constructed from a linear combination of $\mathbf{r}^{\prime}\in{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})$ over ${\mathbb{Z}}_{+}$, so that the condition (3.4) of Lemma 3.2 is satisfied. Now, by virtue of (2.4) and the fact that $\mathbf{b}$ has rational ordinates, it is clear that there exists an integer $\alpha>0$ such that $\alpha\mathbf{b}\in{\mathcal{R}}_{\bm{\gamma}}$. Combining this with the assumption (2.3), we see for all $\mathbf{r}\in{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})$ there exists a natural number $n_{\mathbf{r}}$ such that $n_{\mathbf{r}}\alpha\mathbf{b}-\mathbf{r}\in{\mathcal{R}}_{\bm{\gamma}}$, and thus there exist integers $\lambda_{\mathbf{r}}(\mathbf{r}^{\prime})\geqslant 0$ such that
$n_{\mathbf{r}}\alpha\mathbf{b}-\mathbf{r}=\sum_{\mathbf{r}^{\prime}\in{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})}\lambda_{\mathbf{r}}(\mathbf{r}^{\prime})\mathbf{r}^{\prime}.$
Consequently, we have
$\sum_{\mathbf{r}\in{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})}n_{\mathbf{r}}\alpha\mathbf{b}$ $=\sum_{\mathbf{r}\in{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})}\mathbf{r}+\sum_{\mathbf{r}\in{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})}\sum_{\mathbf{r}^{\prime}\in{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})}\lambda_{\mathbf{r}}(\mathbf{r}^{\prime})\mathbf{r}^{\prime}$
$=\sum_{\mathbf{r}\in{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})}\mathbf{r}\Big{(}1+\sum_{\mathbf{r}^{\prime}\in{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})}\lambda_{\mathbf{r}^{\prime}}(\mathbf{r})\Big{)},$
and thus, letting $\mathbf{r}_{\ell}$ denote the unique element of ${\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})$ associated to the linear form $\ell\in{\mathcal{L}}_{\chi_{\bm{\gamma}}}(\mathbf{a})$, we have
$\sum_{(i,j)\in{\mathcal{I}}^{\mathbf{m}}}b_{i,j}\mathbf{e}_{i,j}^{*}=\sum_{\ell\in{\mathcal{L}}_{\chi_{\bm{\gamma}}}(\mathbf{a})}\beta_{\ell}\ell\quad\text{where}\quad\beta_{\ell}=\Big{(}\sum_{\mathbf{r}\in{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})}n_{\mathbf{r}}\alpha\Big{)}^{-1}\Big{(}1+\sum_{\mathbf{r}\in{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})}\lambda_{\mathbf{r}}(\mathbf{r}_{\ell})\Big{)}>0.$
Therefore, the condition (3.3) of Lemma 3.2 is satisfied. Hence, appealing to the fact that $G_{\psi_{{\mathcal{N}}}\chi_{\bm{\gamma}}}(\mathbf{a},\mathbf{0}_{\mathbf{m}})$ is non-zero, by virtue of Proposition 3.4, we see that all the conditions of Lemma 3.2 are satisfied, and thus the polynomial $P$ satisfies
(4.2) $P(\log H)=\frac{G_{\psi_{{\mathcal{N}}}\chi_{\bm{\gamma}}}(\mathbf{a},\mathbf{0}_{\mathbf{m}})\>{\mathrm{vol\,}}\Omega_{{\mathcal{L}}_{\chi_{\bm{\gamma}}}(\mathbf{a})}(H^{\mathbf{b}})}{H^{\langle\mathbf{a},\mathbf{b}\rangle}}+O\big{(}(\log H)^{\kappa_{\bm{\gamma}}(\mathbf{a})-1}\big{)}$
as $H\to\infty$. Now, for any $\sigma>0$, we have
$G_{\psi_{{\mathcal{N}}}\chi_{\bm{\gamma}}}(\mathbf{a},\sigma\mathbf{1}_{\mathbf{m}})$ $=D_{\psi_{{\mathcal{N}}}\chi_{\bm{\gamma}}}(\mathbf{a}+\sigma\mathbf{1}_{\mathbf{m}})\prod_{\mathbf{r}\in{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})}\sigma\|\mathbf{r}\|$
$=\Big{(}\prod_{\mathbf{r}\in{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})}\sigma\|\mathbf{r}\|\Big{)}\prod_{p}\sum_{n_{1}\geqslant 0}\sum_{n_{2}\geqslant 0}\frac{\nu_{\bm{\gamma},{\mathcal{N}}}(\mathbf{a};n_{1},n_{2},p)}{p^{\sigma n_{1}+n_{2}}}$
(4.3) $=\Big{(}\prod_{\mathbf{r}\in{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})}\sigma\|\mathbf{r}\|\zeta(1+\sigma\|\mathbf{r}\|)\Big{)}\prod_{p}\sum_{n_{1}\geqslant 0}\sum_{n_{2}\geqslant 0}\frac{(1-p^{-1-\sigma\|\mathbf{r}\|})\nu_{\bm{\gamma},{\mathcal{N}}}(\mathbf{a};n_{1},n_{2},p)}{p^{\sigma n_{1}+n_{2}}},$
since, for any $\delta>0$, both $\zeta(1+\delta)$ and $1/\zeta(1+\delta)$ have absolutely convergent Euler products. Here, $\nu_{\bm{\gamma},{\mathcal{N}}}(\mathbf{a};n_{1},n_{2},p)$ denotes the number of $\mathbf{r}\in{\mathcal{R}}_{\bm{\gamma},{\mathcal{N}}}(p)$ with $\|\mathbf{r}\|=n_{1}$ and $\langle\mathbf{a},\mathbf{r}\rangle=n_{2}$. Now, noting that $G_{\psi_{{\mathcal{N}}}\chi_{\bm{\gamma}}}(\mathbf{a},\mathbf{s})$ is, by virtue of Proposition 3.4, holomorphic at $\mathbf{s}=\mathbf{0}_{\mathbf{m}}$, and that $\zeta(s)$ has a simple pole at $s=1$ of residue 1, we derive from (4.2) that
(4.4) $G_{\psi_{{\mathcal{N}}}\chi_{\bm{\gamma}}}(\mathbf{a},\mathbf{0}_{\mathbf{m}})=\prod_{p}(1-p^{-1})^{\#{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})}\sum_{n\geqslant 0}\frac{\nu_{\bm{\gamma},{\mathcal{N}}}(\mathbf{a};n,p)}{p^{n}}.$
[…]
It now remains only to establish an asymptotic for ${\mathrm{vol\,}}\Omega_{{\mathcal{L}}_{\chi_{\bm{\gamma}}}(\mathbf{a})}(H^{\mathbf{b}})$, for then Theorem 2.2 would follow from Proposition 4.2 with ${\mathcal{N}}={\mathbb{N}}^{\mathbf{m}}$ on noting that $\nu_{\bm{\gamma},{\mathbb{N}}^{\mathbf{m}}}(\mathbf{a};n,p)=\nu_{\bm{\gamma}}(\mathbf{a};n)$ for all natural numbers $n$ and primes $p$. Note that expressing the aforementioned volume as a simple formula is unfortunately not possible in full generality. Of course, in any fixed case the volume of $\Omega_{{\mathcal{L}}_{\chi_{\bm{\gamma}}}(\mathbf{a})}(H^{\mathbf{b}})$ can be calculated explicitly, as has been done in the small cases considered in [42, 7, 27]. The best we can achieve in full generality is to relate ${\mathrm{vol\,}}\Omega_{{\mathcal{L}}_{\chi_{\bm{\gamma}}}(\mathbf{a})}(H^{\mathbf{b}})$ to the simpler expression $V_{\bm{\gamma}}(\mathbf{a},\mathbf{b})$ in Theorem 2.1, which the following result achieves.
Lemma 4.3.
Under the hypothesis of Proposition 4.2, we have
$\frac{{\mathrm{vol\,}}\Omega_{{\mathcal{L}}_{\chi_{\bm{\gamma}}}(\mathbf{a})}(H^{\mathbf{b}})}{H^{\langle\mathbf{a},\mathbf{b}\rangle}}\sim V_{\bm{\gamma}}(\mathbf{a},\mathbf{b})(\log H)^{\kappa_{\bm{\gamma}}(\mathbf{a})}\quad\text{as}\quad H\to\infty,$
where $V_{\bm{\gamma}}(\mathbf{a},\mathbf{b})$ is defined in Theorem 2.2.
Proof.
Let $\mathbf{r}_{\ell}$ denote the unique element of ${\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})$ associated to the linear form $\ell\in{\mathcal{L}}_{\chi_{\bm{\gamma}}}(\mathbf{a})$. Write
${\mathcal{T}}_{\mathbf{a},\mathbf{b}}=\Big{\{}(t_{\mathbf{r}})_{\mathbf{r}}\in{\mathbb{R}}_{+}^{{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})}:\sum_{\mathbf{r}\in{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})}r_{i,j}t_{\mathbf{r}}\leqslant b_{i,j}\>\forall(i,j)\in{\mathcal{I}}^{\mathbf{m}}\Big{\}},$
and note that, by the change of variables $y_{\ell}=\exp(t_{\mathbf{r}_{\ell}}\log H)$, we have
(4.5) ${\mathrm{vol\,}}\Omega_{{\mathcal{L}}_{\chi_{\bm{\gamma}}}(\mathbf{a})}(H^{\mathbf{b}})=(\log H)^{\#{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})}\int\limits_{{\mathcal{T}}_{\mathbf{a},\mathbf{b}}}\prod_{\mathbf{r}\in{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})}\exp(t_{\mathbf{r}}\log H)\>\mathrm{d}t_{\mathbf{r}}.$
Now, suppose that ${\mathcal{X}}\subset{\mathbb{R}}_{+}^{\mathbf{m}}$ is a minimal basis for ${\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})$, and for each $\mathbf{r}\in{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})$, define the unique real numbers $\theta_{\mathbf{x},\mathbf{r}}$ to be such that
$\sum_{\mathbf{x}\in{\mathcal{X}}}\theta_{\mathbf{x},\mathbf{r}}\mathbf{x}=\mathbf{r}.$
Without loss of generality, suppose that ${\mathcal{X}}$ is chosen so that $\lambda_{\mathbf{x}}=\theta_{\mathbf{x},\mathbf{b}}>0$ for all $\mathbf{x}\in{\mathcal{X}}$, and such that
$\sum_{\mathbf{x}\in{\mathcal{X}}}x_{i,j}\sigma_{\mathbf{x}}\leqslant b_{i,j}\>\forall(i,j)\in{\mathcal{I}}^{\mathbf{m}}\quad\text{if and only if}\quad\sigma_{\mathbf{x}}\leqslant\lambda_{\mathbf{x}}\>\forall\mathbf{x}\in{\mathcal{X}},$
which clearly can be done. Then, write
${\mathcal{S}}_{\mathbf{a},\mathbf{b}}(H)=\Big{\{}(\sigma_{\mathbf{x}})_{\mathbf{x}}\in{\mathbb{R}}_{+}^{{\mathcal{X}}}:\sigma_{\mathbf{x}}\leqslant\lambda_{\mathbf{x}}\>\forall\mathbf{x}\in{\mathcal{X}},\sum_{(i,j)\in{\mathcal{I}}^{\mathbf{m}}}a_{i,j}\sum_{\mathbf{x}\in{\mathcal{X}}}x_{i,j}\sigma_{\mathbf{x}}\geqslant\langle\mathbf{a},\mathbf{b}\rangle-\frac{1}{\sqrt{\log H}}\Big{\}},$
and, for any $(\sigma_{\mathbf{x}})_{\mathbf{x}}\in{\mathbb{R}}_{+}^{{\mathcal{X}}}$, define
$A(\sigma_{\mathbf{x}})_{\mathbf{x}}={\mathrm{vol\,}}{\mathcal{A}}(\sigma_{\mathbf{x}})_{\mathbf{x}}\quad\text{where}\quad{\mathcal{A}}(\sigma_{\mathbf{x}})_{\mathbf{x}}=\Big{\{}(t_{\mathbf{r}})_{\mathbf{r}}\in{\mathbb{R}}_{+}^{{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})}:\sum_{\mathbf{r}\in{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})}\theta_{\mathbf{x},\mathbf{r}}t_{\mathbf{r}}=\sigma_{\mathbf{x}}\>\forall\mathbf{x}\in{\mathcal{X}}\Big{\}}.$
Note that ${\mathcal{A}}(\sigma_{\mathbf{x}})_{\mathbf{x}}$ is of dimension $\kappa_{\bm{\gamma}}(\mathbf{a})$ whenever $\sigma_{\mathbf{x}}>0$ for all $\mathbf{x}\in{\mathcal{X}}$, by virtue of the fact that $\operatorname{rank}{\mathcal{X}}=\#{\mathcal{X}}=\operatorname{rank}{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})$. It is also clear that there exists a fixed embedding ${\mathbb{R}}_{+}^{\kappa_{\bm{\gamma}}(\mathbf{a})}\cong E\subset{\mathbb{R}}_{+}^{{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})}$ such that ${\mathcal{A}}(\sigma_{\mathbf{x}})_{\mathbf{x}}\subset E$ as $(\sigma_{\mathbf{x}})_{\mathbf{x}}\uparrow(\lambda_{\mathbf{x}})_{\mathbf{x}}$, and thus the conditions of Lemma 3.7 are fulfilled.
Now, we may rewrite the integral on the right-hand side of (4.5) as
$\int\limits_{{\mathcal{T}}_{\mathbf{a},\mathbf{b}}}\prod_{\mathbf{r}\in{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})}\exp(t_{\mathbf{r}}\log H)\>\mathrm{d}t_{\mathbf{r}}$
$\hskip 42.67912pt=\int\limits_{{\mathbb{R}}_{+}^{{\mathcal{X}}}}\exp\Big{(}\log H\sum_{(i,j)\in{\mathcal{I}}^{\mathbf{m}}}a_{i,j}\sum_{\mathbf{x}\in{\mathcal{X}}}x_{i,j}\sigma_{\mathbf{x}}\Big{)}\>A(\sigma_{\mathbf{x}})_{\mathbf{x}}\prod_{\mathbf{x}\in{\mathcal{X}}}\mathrm{d}\sigma_{\mathbf{x}}$
$\hskip 42.67912pt=\int\limits_{{\mathcal{S}}_{\mathbf{a},\mathbf{b}}(H)}\exp\Big{(}\log H\sum_{(i,j)\in{\mathcal{I}}^{\mathbf{m}}}a_{i,j}\sum_{\mathbf{x}\in{\mathcal{X}}}x_{i,j}\sigma_{\mathbf{x}}\Big{)}\>A(\sigma_{\mathbf{x}})_{\mathbf{x}}\prod_{\mathbf{x}\in{\mathcal{X}}}\mathrm{d}\sigma_{\mathbf{x}}+O_{\mathbf{a},\mathbf{b}}(H^{\langle\mathbf{a},\mathbf{b}\rangle}e^{-\sqrt{\log H}})$
(4.6) $\hskip 42.67912pt=(V_{\bm{\gamma}}(\mathbf{a},\mathbf{b})+o_{\mathbf{a},\mathbf{b}}(1))\int\limits_{{\mathcal{S}}_{\mathbf{a},\mathbf{b}}(H)}\prod_{\mathbf{x}\in{\mathcal{X}}}\exp\Big{(}\log H\sum_{(i,j)\in{\mathcal{I}}^{\mathbf{m}}}a_{i,j}x_{i,j}\sigma_{\mathbf{x}}\Big{)}\>\mathrm{d}\sigma_{\mathbf{x}}+O_{\mathbf{a},\mathbf{b}}(H^{\langle\mathbf{a},\mathbf{b}\rangle}e^{-\sqrt{\log H}}),$
by virtue of Lemma 3.7, Hölder’s inequality, and the fact that $A(\lambda_{\mathbf{x}})_{\mathbf{x}}=V_{\bm{\gamma}}(\mathbf{a},\mathbf{b})$. In order that it may become separable, we extend the integral on the right-hand side of (4.2), with negligible error, to derive
$\int\limits_{{\mathcal{S}}_{\mathbf{a},\mathbf{b}}(H)}\prod_{\mathbf{x}\in{\mathcal{X}}}\exp\Big{(}\log H\sum_{(i,j)\in{\mathcal{I}}^{\mathbf{m}}}a_{i,j}x_{i,j}\sigma_{\mathbf{x}}\Big{)}\>\mathrm{d}\sigma_{\mathbf{x}}$
$\hskip 128.0374pt=\prod_{\mathbf{x}\in{\mathcal{X}}}\int\limits_{-\infty}^{\lambda_{\mathbf{x}}}\exp\Big{(}\log H\sum_{(i,j)\in{\mathcal{I}}^{\mathbf{m}}}a_{i,j}x_{i,j}\sigma_{\mathbf{x}}\Big{)}\>\mathrm{d}\sigma_{\mathbf{x}}+O_{\mathbf{a},\mathbf{b}}(H^{\langle\mathbf{a},\mathbf{b}\rangle}e^{-\sqrt{\log H}})$
$\hskip 128.0374pt=\frac{1}{(\log H)^{\#{\mathcal{X}}}}\prod_{\mathbf{x}\in{\mathcal{X}}}\exp\Big{(}\log H\sum_{(i,j)\in{\mathcal{I}}^{\mathbf{m}}}a_{i,j}x_{i,j}\lambda_{\mathbf{x}}\Big{)}+O_{\mathbf{a},\mathbf{b}}(H^{\langle\mathbf{a},\mathbf{b}\rangle}e^{-\sqrt{\log H}})$
$\hskip 128.0374pt=\frac{H^{\langle\mathbf{a},\mathbf{b}\rangle}}{(\log H)^{\#{\mathcal{X}}}}+O_{\mathbf{a},\mathbf{b}}(H^{\langle\mathbf{a},\mathbf{b}\rangle}e^{-\sqrt{\log H}}),$
[…]
Combining Lemma 4.3 with Proposition 4.2, and taking ${\mathcal{N}}={\mathbb{N}}^{\mathbf{m}}$, we complete the demonstration of Theorem 2.2.
4.3. Proof of Theorems 2.3 and 2.4
In order to use Theorem 2.2 to prove the asymptotic formulæ given in §2.2 for the moments of the divisor function $\tau_{m}(n;H^{\mathbf{c}})$, we need only find appropriate choices for the tuple $\mathbf{a}$ to complement the choices of $\bm{\gamma}$ and $\mathbf{b}$ given in (2.5) and (2.6). This procedure is quite straightforward, and we proceed as follows.
Proof of Theorem 2.3..
Recall from (2.5) that it suffices to estimate $\mathsf{E}(\bm{\gamma};H^{\mathbf{b}})$ with
$\bm{\gamma}=\bigoplus_{i=1}^{k}\mathbf{1}_{m}\quad\text{and}\quad\mathbf{b}=\bigoplus_{i=1}^{k}\mathbf{c}.$
Firstly note that, in this case, we have
${\mathcal{R}}_{\bm{\gamma}}=\{r\in{\mathbb{Z}}_{+}^{\mathbf{m}}\colon\|\mathbf{r}_{1}\|=\dots=\|\mathbf{r}_{k}\|\}.$
Moreover, our choice of $\mathbf{b}$ trivially satisfies the condition (2.4), and therefore we need only check that the condition (2.3) is satisfied by the tuple
$\mathbf{a}=\bigoplus_{i=1}^{k}\dfrac{1}{k}\mathbf{1}_{m}\in\bigoplus_{i=1}^{k}{\mathbb{Q}}^{m}.$
In order to prove that this condition is met, firstly note that, for all non-zero $\mathbf{r}\in{\mathcal{R}}_{\bm{\gamma}}$, we must have $\|\mathbf{r}_{i}\|\geqslant 1$ for all $i\leqslant k$, and thus
(4.7) $\langle\mathbf{a},\mathbf{r}\rangle=\frac{1}{k}\|\mathbf{r}\|=\frac{1}{k}(\|\mathbf{r}_{1}\|+\cdots+\|\mathbf{r}_{k}\|)\geqslant 1.$
Moreover, it is clear that equality holds in (4.7) precisely when $\|\mathbf{r}_{i}\|=1$ for all $i\leqslant k$, and consequently
${\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})=\left\{\bigoplus_{i=1}^{k}\bm{\delta}^{m}({w_{i}})\colon w_{i}\in\{1,\ldots,m\}\>\forall i\leqslant k\right\},$
where $\bm{\delta}^{m}(w)\in\{0,1\}^{m}$ is such that $\delta_{j}^{m}(w)=1$ if and only if $j=w$. Hence, it is clear that the condition (2.3) is satisfied by our choice of $\mathbf{a}$, and moreover that $V_{\bm{\gamma}}(\mathbf{a},\mathbf{b})=V_{m,k,\mathbf{c}}$. Thus, appealing to Theorem 2.2, we now need only show that
(4.8) #$\#{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})=m^{k}\quad\text{and}\quad\operatorname{rank}{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})=(m-1)k+1,$
and that, for any natural number $n$, we have
(4.9) $\nu_{\bm{\gamma}}(\mathbf{a};n)=\binom{n+m-1}{m-1}^{\!\!k}.$
The first part of (4.8) is trivial, and we will prove the second part presently by showing that the set
${\mathcal{B}}=\bigcup_{i\leqslant k}\Big{\{}\bigoplus_{t\leqslant k}\bm{\delta}^{m}(w_{t})\colon w_{i}\in\{1,\ldots,m\},w_{t}=1\>\forall t\neq i\Big{\}}$
is linearly independent, and is moreover a basis for $\operatorname{span}{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})$. Firstly, to show that ${\mathcal{B}}$ is linearly independent, it suffices to show that, if $\mu_{\bm{\varrho}}$ are real numbers satisfying
(4.10) $\sum_{\bm{\varrho}\in{\mathcal{B}}}\mu_{\bm{\varrho}}\bm{\varrho}=\bigoplus_{i=1}^{k}\mathbf{0}_{m},$
then we must have $\mu_{\bm{\varrho}}=0$ for all $\bm{\varrho}\in{\mathcal{B}}$. To prove that this is the case, we firstly note that if $\bm{\varrho}=\bm{\delta}^{m}(w_{1})\oplus\cdots\oplus\bm{\delta}^{m}(w_{m})\in{\mathcal{B}}$ is such that $w_{i}\neq 1$ for some $i\leqslant m$, then $\varrho_{i,w_{i}}=1$, while every $\bm{\varrho}^{\prime}\in{\mathcal{B}}\backslash\{\bm{\varrho}\}$ must satisfy $\varrho_{i,w_{i}}^{\prime}=0$, which forces $\mu_{\bm{\varrho}}=0$. Consequently, we must have $\mu_{\bm{\varrho}}=0$ in (4.10) for all $\bm{\varrho}\in{\mathcal{B}}\backslash\{\bm{\delta}^{m}(1)\oplus\cdots\oplus\bm{\delta}^{m}(1)\}$, which is sufficient to prove that ${\mathcal{B}}$ is linearly independent.
Now, to prove that ${\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})$ is contained in the span of ${\mathcal{B}}$, firstly consider the subsets
${\mathcal{R}}_{\bm{\gamma}}(\mathbf{a};u)=\Big{\{}\bigoplus_{t\leqslant k}\bm{\delta}^{m}(w_{i})\in{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a}):\#\{i\leqslant k:w_{i}\neq 1\}=u\Big{\}},$
which clearly satisfy
$\bigcup_{u\leqslant k}{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a};u)={\mathcal{R}}_{\bm{\gamma}}(\mathbf{a}).$
In other words, ${\mathcal{R}}_{\bm{\gamma}}(\mathbf{a};u)$ is the set of vectors in ${\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})$ such that exactly $u$ parts of it are not equal to $\bm{\delta}^{m}(1)=(1,\dots,0)$.
[…]
$\mathbf{r}=\bm{\delta}^{m}(w_{1})\oplus\bm{\delta}^{m}(w_{2})\bigoplus_{i=3}^{k}\bm{\delta}^{m}(w_{i})\quad\text{with}\quad w_{1},\>w_{2}\neq 1.$
In this case, we can rewrite $\mathbf{r}$ as
$\mathbf{r}=\Big{(}\bm{\delta}^{m}(1)\oplus\bm{\delta}^{m}(w_{2})\bigoplus_{i=3}^{k}\bm{\delta}^{m}(w_{i})\Big{)}+\Big{(}\bm{\delta}^{m}(w_{1})\oplus\bm{\delta}^{m}(1)\bigoplus_{i=3}^{k}\bm{\delta}^{m}({w_{i}})\Big{)}-\Big{(}\bm{\delta}^{m}({1})\oplus\bm{\delta}^{m}({1})\bigoplus_{i=3}^{k}\bm{\delta}^{m}({w_{i}})\Big{)},$
where the first and second terms are elements of ${\mathcal{R}}_{\bm{\gamma}}(\mathbf{a};u-1)$, while the last term is an element of ${\mathcal{R}}_{\bm{\gamma}}(\mathbf{a};u-2)$. Consequently, we have
${\mathcal{R}}_{\bm{\gamma}}(\mathbf{a};u)\subset\operatorname{span}\big{(}{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a};u-1),{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a};u-2)\big{)}\subseteq\operatorname{span}{\mathcal{B}},$
the last inclusion following from the induction assumption. This completes the induction step and thus proves the second part of (4.8), and consequently also proves that
(4.11) $\kappa_{\bm{\gamma}}(\mathbf{a})=\#{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})-\operatorname{rank}{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})=m^{k}-(m-1)k-1.$
Now, for any natural number $n$, we note that
$\nu_{\bm{\gamma}}(\mathbf{a};n)=\#\{\mathbf{r}\in{\mathcal{R}}_{\bm{\gamma}}:\langle\mathbf{a},\mathbf{r}\rangle=n\}=\#\{\mathbf{r}\in{\mathcal{R}}_{\bm{\gamma}}:\|\mathbf{r}_{i}\|=n\>\forall i\leqslant k\},$
from which (4.9) follows using a stars-and-bars argument. Hence, appealing to (4.8), (4.9), and (4.11) as well as the fact that
$\langle\mathbf{a},\mathbf{b}\rangle=\frac{1}{k}(\langle\mathbf{1}_{m},\mathbf{c}\rangle+\cdots+\langle\mathbf{1}_{m},\mathbf{c}\rangle)=\|c\|,$
we see by Theorem 2.2 that there exists an $\vartheta>0$ such that
$\mathsf{E}(\bm{\gamma};H^{\mathbf{b}})=H^{\|c\|}Q(\log H)+o(H^{\|c\|-\vartheta}),$
where $Q$ is a polynomial satisfying
$Q(\log H)\sim(\log H)^{m^{k}-(m-1)k-1}V_{\bm{\gamma}}(\mathbf{a},\mathbf{b})\prod_{p}(1-p^{-1})^{m^{k}}\sum_{n\geqslant 0}\frac{1}{p^{n}}\binom{n+m-1}{m-1}^{\!\!k},$
[…]
Proof of Theorem 2.4..
Similarly to as above, we firstly note that
$\sum_{n}\tau_{m}(n^{\ell};H^{\mathbf{c}})^{k}=\mathsf{E}(\bm{\gamma}^{\prime};H^{\mathbf{b}^{\prime}})\quad\text{with}\quad\bm{\gamma}^{\prime}=\bm{\gamma}\oplus(\ell)\quad\text{and}\quad\mathbf{b}^{\prime}=\mathbf{b}\oplus(\|\mathbf{c}\|/\ell),$
where $\bm{\gamma}$ and $\mathbf{b}$ are as in the proof of Theorem 2.3. We now note that $\mathbf{b}^{\prime}$ trivially satisfies the condition (2.4), and that the tuple
$\mathbf{a}^{\prime}=\bigoplus_{i=1}^{k}\dfrac{1}{(k+1)\ell}\mathbf{1}_{m}\oplus\bigg{(}\dfrac{1}{k+1}\bigg{)}$
satisfies the condition (2.3). Indeed, on noting that, if $\mathbf{r}\in{\mathcal{R}}_{\bm{\gamma}^{\prime}}$ is non-zero, then we must necessarily have
$\|\mathbf{r}_{1}\|=\cdots=\|\mathbf{r}_{k}\|=\ell\|\mathbf{r}_{k+1}\|=\ell N,$
for some integer $N\geq 1$. We see that
(4.12) $\frac{1}{(k+1)\ell}(\|\mathbf{r}_{1}\|+\cdots+\|\mathbf{r}_{k}\|)+\frac{1}{k+1}\|\mathbf{r}_{k+1}\|=\frac{k\ell N}{(k+1)\ell}+\frac{N}{k+1}=N\geqslant 1$
for all non-zero $\mathbf{r}\in{\mathcal{R}}_{\bm{\gamma}^{\prime}}$, which shows that our choice of $\mathbf{a}^{\prime}$ satisfies (2.3). Now, noting that equality holds on the right-hand side of (4.12) precisely when $N=1$, and thus if and only if $\|\mathbf{r}_{i}\|=\ell$ for all $i\leqslant k$, we see that
${\mathcal{R}}_{\bm{\gamma}^{\prime}}(\mathbf{a}^{\prime})=\left\{\Big{(}\sum_{t\leqslant\ell}\bm{\varrho}_{t}\Big{)}\oplus(1)\colon\bm{\varrho}_{t}\in{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})\>\forall t\leqslant\ell\right\},$
where ${\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})$ is defined in the proof of Theorem 2.3.
Now, to complete the demonstration, by virtue of Theorems 2.1 and 2.2, it suffices to show that
(4.13) #$\#{\mathcal{R}}_{\bm{\gamma}^{\prime}}(\mathbf{a}^{\prime})=\binom{m+\ell-1}{m-1}^{\!\!k}\quad\text{and}\quad\operatorname{rank}{\mathcal{R}}_{\bm{\gamma}^{\prime}}(\mathbf{a}^{\prime})=\operatorname{rank}{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})=(m-1)k+1,$
and moreover that
(4.14) $\nu_{\bm{\gamma}^{\prime}}(\mathbf{a}^{\prime};n)=\binom{\ell n+m-1}{m-1}^{\!\!k}$
for any natural number $n$. The first part of (4.13) is trivial, and (4.14) follows from a simple stars-and-bars counting argument. In order to prove that the second part of (4.13) holds, we firstly consider the subset of ${\mathcal{R}}_{\bm{\gamma}^{\prime}}(\mathbf{a}^{\prime})$ given by
${\mathcal{S}}\coloneqq\big{\{}\ell{\bm{\xi}}\oplus(1)\colon{\bm{\xi}}\in{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})\big{\}}.$
Now, it is clear that
(4.15) $\operatorname{rank}{\mathcal{R}}_{\bm{\gamma}^{\prime}}(\mathbf{a}^{\prime})\geqslant\operatorname{rank}{\mathcal{S}}=\operatorname{rank}{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})=(m-1)k+1.$
Next, to prove that ${\mathcal{S}}$ spans ${\mathcal{R}}_{\bm{\gamma}^{\prime}}(\mathbf{a}^{\prime})$, we firstly note that any element $\mathbf{r}\in{\mathcal{R}}_{\bm{\gamma}^{\prime}}(\mathbf{a}^{\prime})$ can be represented as $\mathbf{r}=(\bm{\varrho}_{1}+\cdots+\bm{\varrho}_{\ell})\oplus(1)$, where $\bm{\varrho}_{t}\in{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})$ for all $t\leqslant\ell$. We then see that
$\ell\mathbf{r}=\Big{(}\sum_{t\leqslant k}\ell\bm{\varrho}_{t}\Big{)}\oplus(\ell)=\sum_{t\leqslant\ell}\big{(}\ell\bm{\varrho}_{t}\oplus(1)\big{)}\in\operatorname{span}{\mathcal{S}},$
and thus ${\mathcal{R}}_{\bm{\gamma}^{\prime}}(\mathbf{a}^{\prime})\subseteq\operatorname{span}{\mathcal{S}}$. Together with (4.15), this completes the proof of (4.13), and thus we have
$\kappa_{\bm{\gamma}}(\mathbf{a})=\#{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})-\operatorname{rank}{\mathcal{R}}_{\bm{\gamma}}(\mathbf{a})=\binom{m+\ell-1}{m-1}^{\!\!k}-mk+k-1.$
So, appealing additionally to the fact that
$\langle\mathbf{a},\mathbf{b}\rangle=\dfrac{1}{(k+1)\ell}(\langle\mathbf{1}_{m},\mathbf{c}\rangle+\cdots+\langle\mathbf{1}_{m},\mathbf{c}\rangle)+\frac{1}{k+1}(\|\mathbf{c}\|/\ell)=\frac{k\|\mathbf{c}\|}{(k+1)\ell}+\frac{1}{k+1}(\|\mathbf{c}\|/\ell)=\|\mathbf{c}\|/\ell,$
we see by Theorem 2.2 that there exists an $\vartheta>0$ such that
$\mathsf{E}(\bm{\gamma}^{\prime};H^{\mathbf{b}^{\prime}})=H^{\|c\|/\ell}Q(\log H)+o(H^{\|c\|/\ell-\vartheta}),$
where $Q$ is a polynomial satisfying
$Q(\log H)\sim(\log H)^{\binom{m+\ell-1}{m-1}^{\!\!k}-mk+k-1}V_{\bm{\gamma}^{\prime}}(\mathbf{a}^{\prime},\mathbf{b}^{\prime})\prod_{p}(1-p^{-1})^{\binom{m+\ell-1}{m-1}^{\!\!k}}\sum_{n\geqslant 0}\frac{1}{p^{n}}\binom{\ell n+m-1}{m-1}^{\!\!k}$
[…]
5. Further works
5.1. Non-explicit results
For general choices of $\bm{\gamma}$, it may not always be possible to determine explicitly the degree and leading coefficient of the polynomial $Q$ in Theorem 2.1, on account of the choice of $\mathbf{a}$ not satisfying the conditions of Theorem 2.2. Nonetheless, in some of these cases, it still may be possible to obtain an asymptotic for $\mathsf{E}(\bm{\gamma};H^{\mathbf{b}})$ in terms of a polynomial $Q$ of non-explicit degree and leading coefficient. Indeed, we have the following result, which improves Theorem 2.1 under a minor additional condition.
Proposition 5.1.
Suppose that $\lambda>0$ is such that, for all $\eta>0$ there exists a $c_{\eta}>0$ such that
(5.1) $\mathsf{E}(\bm{\gamma};H^{\mathbf{b}})\geqslant c_{\eta}H^{\lambda-\eta}$
for infinitely-many $H$. If there exists an $\mathbf{a}\succ\mathbf{0}_{\mathbf{m}}$ satisfying the hypothesis of Theorem 2.1, such that $\langle\mathbf{a},\mathbf{b}\rangle=\lambda$, then there exists a non-zero polynomial $R$ and a constant $\vartheta>0$ such that
$\mathsf{E}(\bm{\gamma};H^{\mathbf{b}})=H^{\lambda}R(\log H)+o(H^{\lambda-\vartheta})\quad\text{as}\quad H\to\infty.$
The main advantage of the above result is that the polynomial $R$ is necessarily non-zero, as opposed to the polynomial $Q$ in Theorem 2.1 which may be zero. Thus, the additional condition (5.1) of Proposition 5.1 allows us to establish an asymptotic, rather than simply a majorisation. Of course, an upper bound for the degree of $R$ can easily be determined if $\mathbf{a}$ is known explicitly.
As with Theorems 2.3 and 2.4, we may also apply Proposition 5.1 to study certain types of restricted divisor functions. For example, consider a generalisation of (1.13), given by
$\tau_{{\bm{\xi}}}(n;H^{\mathbf{c}})=\#\{\mathbf{d}\in\mathfrak{B}^{\ell}(H^{\mathbf{c}}):d_{1}^{\xi_{1}}\cdots d_{\ell}^{\xi_{\ell}}=n\},$
where ${\bm{\xi}},\mathbf{c}\in{\mathbb{N}}^{\ell}$ for some fixed natural number $\ell$. It is not difficult to see, similarly to as in the proofs of Theorems 2.3 and 2.4, that we have the relation
$\mathsf{E}({\bm{\xi}}\oplus\cdots\oplus{\bm{\xi}};H^{\mathbf{c}\oplus\cdots\oplus\mathbf{c}})=\sum_{n}\tau_{{\bm{\xi}}}(n;H^{\mathbf{c}})^{k}.$
We additionally note that
$\sum_{n}\tau_{{\bm{\xi}}}(n;H^{\mathbf{c}})=\#\mathfrak{B}^{\ell}(H^{\mathbf{c}})\sim H^{\|\mathbf{c}\|}$
as $H\to\infty$, and moreover that
(5.2) $\sum_{n}\tau_{{\bm{\xi}}}(n;H^{\mathbf{c}})^{k}\geqslant(1+o(1))H^{\|\mathbf{c}\|}$
for any natural number $k\geqslant 2$. Now, following the same argument as in the proof of Theorem 2.3, we see that $\mathbf{a}=(1/k)(\mathbf{1}_{m}\oplus\cdots\oplus\mathbf{1}_{m})$ satisfies the hypothesis of Theorem 2.1 for $\bm{\gamma}={\bm{\xi}}\oplus\cdots\oplus{\bm{\xi}}$ and $\mathbf{b}=\mathbf{c}\oplus\cdots\oplus\mathbf{c}$. Consequently, appealing to (5.2) and the fact that $\langle\mathbf{a},\mathbf{b}\rangle=\|\mathbf{c}\|$, we see by Proposition 5.1 that
(5.3) $\sum_{n}\tau_{{\bm{\xi}}}(n;H^{\mathbf{c}})^{k}\sim H^{\|\mathbf{c}\|}R_{{\bm{\xi}},k}(\log H)\quad\text{as}\quad H\to\infty$
for some non-zero polynomial $R_{{\bm{\xi}},k}$ satisfying
$\deg R_{{\bm{\xi}},k}\leqslant\sum_{n}\#\{j\leqslant k:\xi_{j}=n\}^{k}-\#\{\xi_{j}:j\leqslant m\}-\sum_{n}k\max(0,\#\{j\leqslant k:\xi_{j}=n\}-1).$
This follows by adapting an argument used in the proof of Theorem 2.3.
Now, in the particular case where ${\bm{\xi}}=(1,t)$ for some integer $t>1$, we have $\deg R_{(1,t),k}\leqslant 2-2-0=0$, and thus there exists a constant $D_{t,k,\mathbf{c}}>0$ such that
$\sum_{n}\tau_{(1,t)}(n;H^{\mathbf{c}})^{k}\sim D_{t,k,\mathbf{c}}H^{\|\mathbf{c}\|}\quad\text{as}\quad H\to\infty.$
This is an example of where the lower bound (1.6) is precisely of the correct order. Note that applying Theorem 2.2 to improve (5.3) is not directly possible in this case, due to unavailability of a vector $\mathbf{a}$ satisfying the condition (2.3) of Theorem 2.2. The condition (2.3) itself is necessary in order for the condition (3.4) of Lemma 3.2 be fulfilled. This occurrence happens for most choices of $\bm{\gamma}$ that are not of the form $\bm{\gamma}=\mathbf{1}_{\mathbf{m}}$. In addition, for many choices of $\bm{\gamma}$, we have $\min(\|\mathbf{b}_{1}\|,\ldots,\|\mathbf{b}_{k}\|)>\mu_{\bm{\gamma},\mathbf{b}}$, in which case, the trivial bounds (1.4) and (1.6) differ by a factor greater than $O_{\varepsilon}(H^{\varepsilon})$.
5.2. Imposing further restrictions
Suppose that ${\mathcal{N}}\subseteq{\mathbb{N}}^{\mathbf{m}}$ is as in Proposition 3.3, that $G_{f}$ is as in Lemma 3.2, and moreover define $\psi_{{\mathcal{N}}}$ and $\chi_{\bm{\gamma}}$ to be as in the proof of Proposition 4.1. Then, under the hypothesis of Proposition 4.2, we clearly have the relation
(5.4) $\mathsf{E}(\bm{\gamma},{\mathcal{N}};H^{\mathbf{b}})=\frac{G_{\psi_{{\mathcal{N}}}\chi_{\bm{\gamma}}}(\mathbf{a},\mathbf{0}_{\mathbf{m}})}{G_{\chi_{\bm{\gamma}}}(\mathbf{a},\mathbf{0}_{\mathbf{m}})}\mathsf{E}(\bm{\gamma};H^{\mathbf{b}})+H^{\langle\mathbf{a},\mathbf{b}\rangle}T(\log H)+o(H^{\langle\mathbf{a},\mathbf{b}\rangle-\vartheta})\quad\text{as}\quad H\to\infty$
for some polynomial $T$ of degree not exceeding $\kappa_{\bm{\gamma}}(\mathbf{a})-1$. For example, taking ${\mathcal{N}}$ to be the set of four-tuples of square-free numbers and appealing to (5.8), we see that the number of singular $2\times 2$ matrices with positive square-free entries not exceeding $H$ is asymptotic to
$\frac{2}{\zeta(2)}\prod_{p}\frac{(p-1)^{3}(p^{2}+4p+1)}{p^{4}(p+1)}H^{2}\log H,$
by virtue of (5.4) and the fact that
$\frac{G_{\psi_{{\mathcal{M}}^{4}}\chi_{\bm{\gamma}}}(\tfrac{1}{2}\mathbf{1}_{(2,2)},\mathbf{0}_{(2,2)})}{G_{\chi_{\bm{\gamma}}}(\tfrac{1}{2}\mathbf{1}_{(2,2)},\mathbf{0}_{(2,2)})}=\prod_{p}\frac{(p-1)^{3}}{p^{2}(p+1)}\frac{p^{2}+4p+1}{p^{2}},$
where ${\mathcal{M}}\subset{\mathbb{N}}$ denotes the set of square-free natural numbers. Similarly, we can use (5.8) and (5.4) to show that, for any natural number $q$, we have
$\sum_{\begin{subarray}{c}n\leqslant H^{2}\\ (n,q)=1\end{subarray}}\tau(n;H)^{2}\sim\frac{2}{\zeta(2)}\prod_{p|q}\frac{(p-1)^{3}}{p^{2}(p+1)}H^{2}\log H\quad\text{as}\quad H\to\infty,$
with an error term similar to that in (5.8).
More generally, we can use (5.4) along with (4.4) to count the number of integer solutions to the multiplicative Diophantine equation (1.2) where only select ordinates are restricted. We simply offer the above two examples as a way of illustrating to the reader how this procedure can easily be followed by appealing to Theorem 2.2.
We note that our method does not, unfortunately, allow us to count solutions to (1.2) over the Piatetski-Shapiro or Beatty numbers, as in these cases ${\mathcal{N}}$ is not closed under the Hadamard product. This is also the main obstacle for considering the case where we restrict to solutions lying in fixed residue classes. A possible exception of this case is when the class of residues are closed under Hadamard product, such as when we consider the set of quadratic or cubic residues modulo $\mathbf{q}$ for some integer vector $\mathbf{q}$, or when all solutions are $1$ modulo $\mathbf{q}$.
5.3. Other future directions
We note that our method can be adapted to count the number of solutions $\mathbf{x}\in\mathfrak{B}^{\mathbf{m}}(H^{\mathbf{b}})$ to equations of the form
(5.5) $\alpha_{1}\mathbf{x}_{1}^{\bm{\gamma}_{1}}=\cdots=\alpha_{k}\mathbf{x}_{k}^{\bm{\gamma}_{k}},$
where $\bm{\alpha}\in{\mathbb{N}}^{k}$. While the corresponding function $f$ for use in Lemmata 3.1 and 3.2 is clearly not multiplicative, it is possible to split the problem into a finite number of counting problems of the form
[…]
where $U_{i,j}=H^{b_{i,j}}/h_{i,j}$ for some small integer solution $\mathbf{h}$ to (5.5). Even more generally, we could consider the problem of counting $\mathsf{E}(\bm{\gamma};\mathbf{X},\mathbf{Y})$, the number of integer solutions to (1.2) lying in the parallelpiped
${\mathcal{P}}^{\mathbf{m}}(\mathbf{X},\mathbf{Y})=\big{\{}\mathbf{d}\in{\mathbb{Z}}^{\mathbf{m}}:X_{i,j}<d_{i,j}\leqslant X_{i,j}+Y_{i,j}\>\forall(i,j)\in{\mathcal{I}}^{\mathbf{m}}\big{\}},$
where $\mathbf{X},\mathbf{Y}\succ\mathbf{0}_{\mathbf{m}}$. In particular, it would be interesting to determine the behaviour of $\mathsf{E}(\bm{\gamma};\mathbf{X},\mathbf{Y})$ as each $X_{i}$ and $Y_{i}$ independently tend to infinity.
Yet another adaptation of our method can be used to count the number of solutions to equations of the form
$\alpha_{1}\mathbf{x}_{1}^{\bm{\gamma}_{1}}=\cdots=\alpha_{k}\mathbf{x}_{k}^{\bm{\gamma}_{k}}\pmod{K},$
for a fixed $K$, in a parallelpiped centered at the origin. The simplest case, $x_{1}x_{2}\equiv x_{3}x_{4}\pmod{p}$ when $k=p$ is a prime, was studied explicitly by Ayyad, Cochrane and Zheng [3], Garaev and Garcia [24] and Kerr [30] in an arbitrary box $B$ with character sums. We note that their method gives the best current result for counting solutions to the equation $x_{1}x_{2}=x_{3}x_{4}$ in a box, as stated in (1.12).
Another interesting direction for the case $k=2$ is to consider the statistics of the integer equation
(5.6) $x_{1}\dots x_{m}-y_{1}\dots y_{m}=\Delta,$
for $m\geq 2$, $\Delta\neq 0$. For the case $m=2$, this problem corresponds to counting the number of $2\times 2$ integer matrices with determinant $\Delta$. Our method can not be directly applied to study the equation (5.6), due to the fact that the corresponding solution set is not closed under the Hadamard product. However, using different methods, the first-named author [1] and Ganguly and Guria [21] have, in the case $m=2$, given asymptotic formulæ for the number of solutions to (5.6) in the box $[-H,H]^{4}$, as $H$ and $|\Delta|$ both tend to infinity.
Motivated by the applications of such results in the works [41, 38], Pappalardi, Sha, Shparlinski, and Stewart [39] studied the multiplicative dependency of $n$-tuples of algebraic integers of bounded height. Related to our work is their study of, essentially, the number
[…]
of multiplicatively dependent positive integer $n$-tuples. In particular, their work can be used to compute the constant $C_{n}>0$ with the property that
(5.7) $\mathsf{M}_{n}(H)=C_{n}H^{n-1}+O\big{(}H^{n-2+\vartheta_{n}/\log\log H}\big{)}\quad\text{as}\quad H\to\infty$
for some constant $\vartheta_{n}>0$. This work leads us to consider generalisations of our own work, such as the problem of determining the asymptotic behaviour of the number
[…]
as $H$ tends to infinity, where ${\mathcal{C}}\subset{\mathbb{N}}^{\mathbf{m}}$ is fixed. Now, the proof of (5.7) given in [39] does not appeal to the work of de la Bretèche [8, 9], but rather makes critical use of a result of Widmer [48]. This is due to the fact that they actually prove a much more general result than (5.7), wherein the sum $S_{f}(X^{\mathbf{b}})$ is not relatable to the number of interest. Nonetheless, it is likely that Lemmata 3.1 and 3.2 of the present article could be used to derive an asymptotic estimate for $\mathsf{E}({\mathcal{C}};H^{\mathbf{b}})$ in the case where ${\mathcal{C}}$ is sufficiently well-structured.
5.4. Applications to counting singular matrices
Consider the sum of interest in Theorem 2.3 in the case where $m=k=2$ and $\mathbf{c}=\mathbf{1}_{2}$. Note that $V_{2,2,\mathbf{1}_{2}}$ is the $12^{2}-2(2-1)-1=$ 1 dimensional volume of a region embedded in a $42^{2}=$ 4 dimensional space, that is, the length of a line segment in ${\mathbb{R}}^{4}$. Therefore
$V_{2,2,\mathbf{1}_{2}}$ $={\mathrm{vol\,}}\{\textbf{v}\in{\mathbb{R}}_{+}^{4}:v_{1}+v_{3}=v_{1}+v_{4}=v_{2}+v_{3}=v_{2}+v_{4}=1\}=\sqrt{1^{2}+1^{2}+1^{2}+1^{2}}=2,$
and consequently
$q_{2,2,\mathbf{1}_{2}}=2\prod_{p}(1-p^{-1})^{4}\sum_{n\geqslant 0}\frac{(n+1)^{2}}{p^{n}}=2\prod_{p}(1-p^{-2})=\frac{2}{\zeta(2)}.$
This allows us to show that there exists a constant $\vartheta>0$ such that
(5.8) $\sum_{n}\tau(n;H)^{2}=\frac{2}{\zeta(2)}H^{2}\log H+AH^{2}+o(H^{2-\vartheta})$
for some constant $A$. Of course, the constant $A$ is the same as the one given in (1.12), which supersedes (5.8) in that the power-saving is explicit. Indeed, for the smaller cases such as this, elementary methods tend to give stronger results than the more general machinery of de la Bretèche, which is best-suited to the larger systems of variables considered in this article.
We now note that (5.8), and thus (1.12), improve on a key lemma of the first-named author [1] used for counting the number of elements in the set ${\mathcal{D}}_{2}(H;\Delta)$ of $2\times 2$ integer matrices of determinate $\Delta$ whose entries are integers of absolute value not exceeding $H$. Asymptotic formulæ for #$\#{\mathcal{D}}_{2}(H;\Delta)$ when $\Delta\neq 0$ have been given in [1, 21, 25] for a large range of $\Delta$ relative to $H$. For $\Delta=0$, the first-named author [1] showed that
#$\#{\mathcal{D}}_{2}(H;0)=\dfrac{96}{\pi^{2}}H^{2}\log H+O(H^{2}).$
We may now use (1.12) to improve this to the following.
Proposition 5.2.
As $H\to\infty$, we have
#$\#{\mathcal{D}}_{2}(H;0)=\dfrac{96}{\pi^{2}}H^{2}\log H+(8A+16)H^{2}+O(H^{19/13}(\log H)^{7/13}),$
where $A$ is as in (1.12).
Proof.
Using the determinant formula, we see that #$\#{\mathcal{D}}_{2}(H;0)$ is the number of integer solutions $(a,b,c,d)$ to the equation
(5.9) $ad=bc,$
with $-H\leq a,b,c,d\leq H$.
We first consider the case where at least one of $a$, $b$, $c$, and $d$ is zero. In fact, we may see that at least one element of $\{a,d\}$ and $\{b,c\}$ is zero. When $a=0$, we have $12H+$ 1 choices of $d$ and $14H+$ 1 choices for $(b,c)$, since one of $b$ and $c$ is zero. Next, when $a\neq 0$ (with $2H$ choices), we have $d=0$ and we again have $14H+$ 1 choices for $(b,c)$. Hence, this case gives
$(4H+1)^{2}=16H^{2}+8H+1$
solutions to (5.9).
We may now assume $abcd\neq 0$. By considering all possible signs of $a$, $b$, $c$ and $d$, similar to as in §3 of [1], we see that
#$\#{\mathcal{D}}_{2}(H;0)=8\sum_{n}\tau(n;H)^{2}+16H^{2}+O(H),$
[…]
Another consequence of (1.12) concerns the problem of counting the number of $2\times 2$ singular matrices whose entries are Egyptian fractions (or unit fractions) of a bounded height, first introduced in [2]. Following the notation in [2], we consider the problem of counting the cardinality of
${\mathcal{D}}_{2}({\mathbb{Z}}^{-1};H,0)=\Big{\{}(a,b,c,d)\in{\mathbb{Z}}^{4}\colon\det\begin{pmatrix}1/a&1/b\\ 1/c&1/d\end{pmatrix}=0,1\leqslant|a|,|b|,|c|,|d|\leqslant H\Big{\}}.$
The first-named author, along with Kuperberg, Ostafe and Shparlinski [2], proved that
$H^{2}\ll\#{\mathcal{D}}_{2}({\mathbb{Z}}^{-1};H,0)\mathop{\>\ll\>}_{\varepsilon}H^{2+\varepsilon}.$
By employing a nearly identical argument as in the proof of Proposition 5.2, we have the following asymptotical formula for #$\#{\mathcal{D}}_{2}({\mathbb{Z}}^{-1};H,0)$.
Proposition 5.3.
As $H\to\infty$, we have
#$\#{\mathcal{D}}_{2}({\mathbb{Z}}^{-1};H,0)=\dfrac{96}{\pi^{2}}H^{2}\log H+8AH^{2}+O(H^{19/13}(\log H)^{7/13}),$
where $A$ is as in (1.12).
Acknowledgements
The authors would like to thank Alina Ostafe, Igor E. Shparlinski, and Liangyi Zhao for supports and suggestions during the preparation of this work. The authors would also like to thank Régis de la Bretèche, as well as Marc Munsch and Igor E. Shparlinski, for helpful correspondences regarding their articles [8], and [36], respectively. The authors would again like to thank Marc Munsch for bringing the articles [31, 3] to their attention. M.A. is supported by Australian Research Council Grants DP230100530 and a UNSW Tuition Fee Scholarship. C.C.C. is supported by the Australian Government and UNSW through RTP and UPA Scholarships, respectively.